### v1

# UniHENN: Designing More Versatile Homomorphic Encryption-based CNNs without im2col

Abstract
Homomorphic encryption enables computations on encrypted data without decryption, which is crucial for privacy-preserving cloud services. However, deploying convolutional neural networks (CNNs) with homomorphic encryption encounters significant challenges, particularly in converting input data into a two-dimensional matrix for convolution, typically achieved using the im2col technique. While efficient, this method limits the variety of deployable CNN models due to compatibility constraints with the encrypted data structure. UniHENN, a homomorphic encryption-based CNN architecture, eliminates the need for im2col, ensuring compatibility with a diverse range of CNN models using homomorphic encryption. Our experiments demonstrate that UniHENN surpasses the leading 2D CNN inference architecture, PyCrCNN, in inference time, as evidenced by its performance on the LeNet-1 dataset, where it averages 30.090 seconds–significantly faster than PyCrCNN’s 794.064 seconds. Furthermore, UniHENN outperforms TenSEAL, which employs im2col, in processing concurrent images, an essential feature for high-demand cloud applications. The versatility of UniHENN is proven across various CNN architectures, including 1D and six different 2D CNNs, highlighting its flexibility and efficiency. These qualities establish UniHENN as a promising solution for privacy-preserving, cloud-based CNN services, addressing the increasing demand for scalable, secure, and efficient deep learning in cloud computing environments.
Index Terms:
homomorphic encryption, privacy-preserving machine learning, convolutional neural networks, data privacy
I Introduction
The growing use of machine learning (ML) in diverse domains such as finance, autonomous driving, industrial automation, computer vision, speech recognition, and healthcare has raised concerns regarding data privacy and security [35, 50]. These concerns are particularly relevant for cloud-based ML services, where sensitive data is processed on remote servers [3].
Homomorphic encryption (HE) is a form of encryption that allows some computations on ciphertext without decrypting it. Therefore, it can be used as a powerful tool for preserving privacy in ML applications [22, 23, 20, 33, 13, 24, 30, 31, 11], enabling the confidentiality of sensitive information even when processed by a third party. HE is particularly relevant for cloud-based ML services, where sensitive data is often transmitted to remote servers for processing. With HE, cloud providers can ensure data privacy in their possession. For instance, a bank could employ HE to build an ML model for detecting fraudulent transactions without accessing customer data. Similarly, a healthcare provider could analyze patients’ medical records using HE without disclosing the patient’s identity. This potential has attracted interest from cloud service providers (e.g., Amazon AWS, Microsoft Azure, Naver Cloud) in integrating HE into ML-based cloud services, such as Data Box Frame [41].
While convolution operations are fundamental to convolutional neural networks (CNNs), they demand substantial computational resources when performed on ciphertexts within an HE framework. Each kernel shift requires element-wise multiplications and additions, resource-intensive operations on ciphertexts. To reduce the computational load, most HE-based CNN implementations like TenSEAL [7] use the im2col function. This function transforms the input data into a matrix without discarding or altering the original information, thus allowing for more efficient computation. TenSEAL is applied in various research areas, including healthcare [21], federated learning [46], and biometrics [49]. However, this approach limits the versatility of CNN models to configurations that accommodate only a single convolution layer, potentially hindering their use in more complex structures. In practical scenarios, encrypted user data could be useful across multiple ML models. For instance, an encrypted patient image stored in a hospital database could be utilized for future analysis by various ML algorithms. Therefore, encrypting data without restricting it to a specific model would offer greater flexibility and utility.
In this paper, we introduce UniHENN, a privacy-preserving CNN model built using HE. UniHENN enables efficient CNN inference independent of the im2col function, resulting in encrypted input data that is usable across other ML models. We devise an effective computation method for convolution operations on flattened input within UniHENN, offering flexible reuse across various ML models. Furthermore, UniHENN supports batch operations, thereby facilitating the efficient processing of multiple ciphertexts. Our key contributions are summarized as follows:
  • We introduce UniHENN, a novel and efficient CNN model inference mechanism based on HE, which facilitates input ciphertext reusability. Unlike other CNN implementations that require a specific model input structure for the im2col function, UniHENN is designed to handle model-free input ciphertexts, enabling the use of encrypted input data across various HE-based ML services without the need for re-encryption. We demonstrate the effectiveness of this approach by successfully constructing and training seven different CNN models on four datasets: MNIST [15], CIFAR-10 [26], USPS [19], and electrocardiogram (ECG) [38]. The source code for UniHENN is available at https://github.com/hm-choi/uni-henn.
  • We empirically demonstrate that UniHENN is efficient. Experimental results indicate an average inference time of 30.090 seconds, significantly outperforming the state-of-the-art HE-based 2D CNN inference architecture PyCrCNN [16], which requires an average of 794.064 seconds for inference.
  • We introduce a batch processing strategy for UniHENN to handle multiple data instances in a single operation. This strategy efficiently combines multiple data instances into a single ciphertext, reducing the inference time for 10 MNIST images to 16.248 seconds. It outperforms TenSEAL’s CNN model [7], which takes 63.706 seconds. While UniHENN is less efficient than TenSEAL for processing a single image, it surpasses TenSEAL’s performance when processing $k$ images simultaneously, where $k\geq 3$.
II Background
II-A Convolutional Neural Network (CNN)
Convolutional neural networks (CNNs) were first introduced by LeCun et al. [28] for processing grid-structured data, such as images. The LeNet-1 model, presented in 1989 [27], comprised two convolutional layers, two average pooling layers, and one fully connected layer, marking the first concept of the LeNet architecture.
In 1998, the LeNet-5 model was introduced [29], consisting of three convolutional layers, two pooling layers, and two fully connected layers. It was designed to recognize handwritten postal codes efficiently, outperforming traditional methods like the multilayer perceptron model.
CNNs gained significant traction in the 2010s, primarily due to their exceptional performance on large datasets like ImageNet [14], which includes millions of images across 1,000 classes. Currently, research into the utilization of CNNs, such as DGCNN and its applications, is ongoing [47].
A typical CNN includes an input layer, multiple hidden layers, and an output layer. The input layer receives the image and forwards its pixel values into the network. Since images are generally matrices, the node count in the input layer corresponds to the image size. The hidden layers, situated between the input and output layers, extract relevant features using convolution, activation, and pooling operations. Each hidden layer processes information from the input or preceding layers, enabling iterative learning. This information is then relayed to subsequent layers, culminating in the final prediction or classification at the output layer.
Kiranyaz et al. [25] introduced a 1D CNN for disease-specific ECG classification. A 1D CNN, tailored for one-dimensional data, is particularly suitable for signal data, time-series data, and text data. In this architecture, both the input and the convolution filter are one-dimensional; the filter slides over the input in the convolution layer to perform operations.
II-B Homomorphic Encryption (HE)
Homomorphic encryption (HE) is a cryptographic technique that allows computations on encrypted data without the need for decryption. Formally, given messages $m_{1}$ and $m_{2}$, an encryption function $Enc$, and computationally feasible functions $f$ and $f^{\prime}$ for ciphertext and plaintext, respectively, HE satisfies $f(Enc(m_{1}),Enc(m_{2}))=Enc(f^{\prime}(m_{1},m_{2}))$.
Several HE schemes, such as BGV [37], GSW-like schemes [8, 10], and CKKS [9], exist, each with different computational needs and data types. In UniHENN, we use the CKKS scheme, which is advantageous for encrypting vectors of real or complex numbers.
Before encrypting with CKKS, it is crucial to create a plaintext structure that mimics the input data. The number of slots specifies the vector size that can be encrypted and is set during parameter selection. CKKS supports three basic operations: addition, multiplication of ciphertexts, and rotation of ciphertext.
Let $N$ be the total degree of the CKKS parameter then the number of slots is $N/2$. Two real vectors $\mathbf{v_{1}},\mathbf{v_{2}}$ and $C(\mathbf{v_{1}}),C(\mathbf{v_{2}})$ denotes the ciphertext of the vector $\mathbf{v_{1}},\mathbf{v_{2}}$. (P) means that the operation is defined between a ciphertext and a plaintext, and (C) means that the operation is defined between two ciphertexts. Then the operations, addition ($Add$), multiplication ($Mul$), and rotation $(Rot)$ can be represented as follows:
  • Addition (P): $Add(C(\mathbf{v_{1}}),\mathbf{v_{2}})=C(\mathbf{v_{1}}+\mathbf{v_{2}})$
  • Addition (C): $Add(C(\mathbf{v_{1}}),C(\mathbf{v_{2}}))=C(\mathbf{v_{1}}+\mathbf{v_{2}})$
  • Multiplication (P): $Mul(C(\mathbf{v_{1}}),\mathbf{v_{2}})=C(\mathbf{v_{1}}\times\mathbf{v_{2}})$
  • Multiplication (C): $Mul(C(\mathbf{v_{1}}),C(\mathbf{v_{2}}))=C(\mathbf{v_{1}}\times\mathbf{v_{2}})$
  • Rotation :
    $Rot(C(\mathbf{v}),r)=C(v_{r},v_{r+1},...,v_{N/2-1},v_{0},...,v_{r-1})$
    where $\mathbf{v}=(v_{1},v_{2},...,v_{N/2})$ and $r$ is a positive integer.
The notation +,×+$\times$+ , × in the plaintext addition and multiplication are elementwisely defined. The depth specifies the maximum number of multiplications, which affects the noise level in the ciphertext, and exceeding this limit may compromise decryption. The level specifies the allowed number of the multiplication. The depth is fixed when the parameter is determined but the level can be decreased when multiplications are conducted.
For practical HE implementations, various libraries are available, including SEAL-Python [2], Lattigo [39], HElib [18], and OpenFHE [5]. We selected SEAL-Python for its efficient support of the CKKS scheme.
II-C HE Libraries
Several libraries have been developed in the HE field to facilitate practical implementations. Among these, SEAL-Python [44], Lattigo [39], HElib [1], and OpenFHE [5] are popularly used.
SEAL [44] supports various HE schemes (BFV and CKKS) and incorporates optimizations to speed up homomorphic operations, such as ciphertext packing and batching. For UniHENN, we chose SEAL-Python, a Python adaptation of the original C++ SEAL library. We selected SEAL-Python for its user-friendly interface. However, UniHENN could also be implemented with other libraries.
Lattigo [39], written in Go, focuses on ring-learning-with-errors-based HE primitives and multiparty-HE-based secure protocols. Its cross-platform compatibility and optimized protocols enable fast and secure computation.
HElib [1] specializes in the BGV and CKKS schemes, incorporating bootstrapping and several optimizations to speed up homomorphic evaluation. It effectively uses the Smart-Vercauteren ciphertext packing techniques and the Gentry-Halevi-Smart optimization. Also, it provides an assembly language for HE, which includes low-level routines, multithreading capabilities, and more.
OpenFHE [5], supporting both bootstrapping and various hardware acceleration backends, utilizes a standard hardware abstraction layer to ensure compatibility and efficiency.
III Overview of UniHENN
In this section, we introduce UniHENN, a HE-based framework designed for inference on encrypted data. UniHENN uses the CKKS HE scheme to encrypt input data, facilitating its integration into any CNN models without needing the im2col function, which demands a specific input shape. To minimize the computational overhead of HE-based inference, we employ three innovative techniques:
  1. 1.
    Unlike previous approaches [7, 16, 18] that rely on the input size, UniHENN calculates the total number of HE operations for the fully connected layer based on the output size. As the output size is typically smaller than the input size in the fully connected layer, this method reduces the average time per operation.
  2. 2.
    UniHENN enables batch operations, allowing the consolidation of multiple ciphertexts into one for more efficient processing. This considerably reduces the overall computational time, marking a significant advantage for UniHENN in large-scale data processing scenarios.
  3. 3.
    UniHENN uses average pooling, eliminating multiplication. This configuration allows for more layers within given parameter settings, giving service providers more flexibility to incorporate average pooling without concern for operation count.
Figure 1: Overview of the UniHENN architecture.
Figure 1 presents a high-level overview of UniHENN’s operational flow. The process begins when a client encrypts one or more images using a public key. These encrypted images are consolidated into a single ciphertext and sent to a cloud service specializing in data analytics. Each service (i.e., Service 1, Service 2, …, Service N) uses its unique CNN model for data processing, executing specific algorithms and computations on the received ciphertext with the evaluation key. Each CNN model has a distinct layer architecture and optimized parameters. After processing the ciphertext through their respective CNN models, the encrypted inference results are returned to the client. The client decrypts these results using the corresponding secret key to obtain the processed outcomes from each CNN model. This framework allows the client to benefit from multiple CNN models while preserving the confidentiality of the data, indicating a significant step forward in privacy-preserving machine learning.
In Table I, we summarize the important notations used in this paper with UniHENN.
TABLE I: Notations used for UniHENN.
Notation
Definition
$N$
Number of slots of ciphertext in power of 2.
$D$
Depth (or level) of ciphertext, representing the current limit on the number of multiplication operations allowed for the ciphertext.
$CH_{in},CH_{out}$
Number of input and output channels in a layer.
$DAT_{in},DAT_{out}$
Number of input and output data in a fully connected (FC) layer.
$R_{q}$
Polynomial quotient ring $(\mathbb{Z}_{q}/\langle X^{N}+1\rangle)$, used to create a vector space for the ciphertext.
$W_{img},H_{img}$
Width and height of the used image data.
$K$
Width and height of the kernel.
$W_{in},H_{in}$
Width and height of input data for this layer.
$W_{out},H_{out}$
Width and height of output data for this layer.
$S_{total}$
Used in Section IV-E. The value is the product of each convolutional layer’s stride and the kernel sizes of all average pooling layers.
III-A Construction of Input Data
There is a variety of data types for ML services. For instance, while image data is two or three-dimensional, statistical data is usually one or two-dimensional. To handle such varying data dimensions, UniHENN initially flattens the input data into a one-dimensional array in row-major format. When performing encryption, the data is located from the first of the list, and the remaining space is filled with zero to perform encryption. The remaining space will be used in the batch operations.
Figure 2 illustrates an example of data transformation for encryption. The input data consists of $W_{img}\times H_{img}$ numbers. We flatten the original data row by row, as shown in Figure 2. Implementing this flattening procedure ensures universal compatibility of the input data with ML models, enhancing its adaptability to the algorithms used by ML services.
Figure 2: Data transform for UniHENN.
III-B Combining Ciphertexts for Batch Operation
The technique in UniHENN combines multiple encrypted input data into a single ciphertext for batch processing. Incorporating multiple encrypted data into a single ciphertext involves sequentially inserting each encrypted data, ensuring enough space to prevent overlap with other encrypted data, as shown in Figure 3. This is achieved by rotating each encrypted data before adding it to the ciphertext.
Importantly, the size allocation is not solely based on the input size of the encrypted data but also considers the size of the intermediate or final output vector generated by performing CNN operations. By examining the given CNN model structure, we can pre-calculate this size in advance, thereby preventing the overlap of input data results.
For example, as illustrated in Figure 3, if 800 slots are needed to process the input data through the CNN model, then the $i$-th input data would be rotated to the right by $(i-1)\times 800$ slots, and the rotated vector is added to the ciphertext. UniHENN can integrate any number of ciphertexts into a single one, provided the total size of the encrypted data does not exceed the number of slots. This method enhances the efficiency of UniHENN when handling multiple ciphertexts concurrently.
Figure 3: Combining multiple ciphertexts into one.
IV CNN Model Construction
IV-A Construction of the Drop-DEPTH
The parameter is pre-defined because UniHENN’s design philosophy is to support the ciphertext that is encrypted before the model is fixed. So, the CKKS parameter is determined and encrypted before the CNN model is defined for reusability of the input ciphertext independent of the model. But, the multiplication and rotation time of HE is linearly increased dependent on the depth. Thus, the best choice of the CKKS parameter is determined by the number and structure of the CNN model’s layers. However, in this scenario, the depth is defined before when the CNN model is fixed, which means that the depth can be larger than the exact number of total levels of the CNN’s model. Thus, we suggest the Drop-DEPTH, a novel approach to reduce the depth that is optimized when the target CNN model. If the input ciphertext has depth $D$ and the target CNN model’s level is $L$ where $D\leq L$ then the Drop-DEPTH reduces the depth of the input ciphertext as $L$ with multiply $D-L$ ciphertexts that are the encryption of vectors where all elements are consists of $1$. By these simple structure, time of execution is very fast (In our experiments in Section VII, the time of Drop-DEPTH operation does not exceed $100ms$.
IV-B Construction of the Convolutional Layer
The im2col encoding is a highly efficient algorithm for converting multidimensional data into matrix form to facilitate convolution operations. However, it requires precise arrangement of input matrix elements, which is challenging to achieve on ciphertexts. Therefore, TenSEAL [7], a sophisticated open-source HE-enabled ML SDK, performs this operation in a preprocessing step before encryption, allowing only a single convolution.
To overcome the limitations of im2col, we introduce a new method for computing convolutional layers that does not rely on this encoding. Our approach uses flattened input data, eliminating the need for a specific arrangement of input matrix elements and thereby increasing flexibility.
In convolutional layers, a kernel traverses an input image with each movement termed a stride. The kernel executes a convolution operation at each stride, multiplying its values with corresponding input image values and summing the results. The input image dimensions are denoted as $(W_{img},H_{img})$, and the stride as $S$. And for simplicity, we define the width and height are equal and denoted as $K$. The total number of kernel movements is given by Equation 1:
$\left(\frac{W_{img}-S}{K}\right)\left(\frac{H_{img}-S}{K}\right)$ (1)
The total number of operations executed by a convolutional layer is influenced by both the size of the input image and the kernel, as indicated by Equation 1. However, given that the input image is usually significantly larger than the kernel, we can approximate that the total number of operations primarily depends on the kernel size.
Figure 4: Convolution operations on the two-dimensional input data and kernel.
As shown in Figure 4, the input data dimension for a basic convolution is two-dimensional. However, input data must be restricted to one-dimensional form in the context of HE. The input data is flattened, as demonstrated in Figure 5, using the method outlined in Section III-A. For simplicity, we assume that both the input data and the kernel are two-dimensional squares and disregard bias operations.
Figure 5: Convolution operations on the flattened input data.
In Figure 5, the number of rotation operations equals the number of elements in the kernel. The input data is incrementally increased by the original kernel size $K$, followed by a leftward rotation operation after each increment. After adding $K\times W_{img}$ cipertexts, another leftward rotation is performed, incrementing by one at each step. The detailed algorithm of these mechanism is described in Algorithm 1. All rotated ciphertexts are then multiplied by the kernel vector, which is specially structured to contain a single kernel element with all other positions set to zero. This eliminates unrelated elements in the ciphertext during convolution, as depicted in Figure 6. For 1D convolution, the logic remains the same, except that the one-dimensional data consists of a single row. The output is an array spaced by the stride interval, as shown in Figure 7.
Figure 6: Flattening a two-dimensional kernel into a one-dimensional array.
Figure 7: Output of convolution operations on one-dimensional data.
This approach has the advantage of decoupling the number of operations from the size of the image. Instead, it performs convolutional operations on transformed two-dimensional data and limits rotations to the square of the kernel size $K$.
Previous approaches have used im2col encoding [7] or performed operations on individual data values [16]. However, the im2col method restricts the model to a single convolutional layer and requires model-specific encoding. Methods that operate on each data value are considerably slower.
Algorithm 1 Construction of the convolution layer
Input:
- List of Ciphertext :
$C_{in}:=(C_{in(0)},C_{in(1)},\dots,C_{in(CH_{in}-1)})\in(\mathcal{R}_{q}^{2})^{CH_{in}}$
- Kernel : $Ker:=(Ker_{(o,i)})_{0\leq o<CH_{out},0\leq i<CH_{in}}$
        $\in\mathbb{R}^{CH_{out}\times CH_{in}\times K\times K}$
- Bias : $B:=(B_{(o)})_{0\leq o<CH_{out}}\in\mathbb{R}^{CH_{out}}$
- Interval : $I_{in}$
- Stride : $S$
Output:
- List of Ciphertexts : $C_{out}\in(\mathcal{R}_{q}^{2})^{CH_{out}}$
- Interval : $I_{out}$
Procedure: *(All ciphertexts and operation are in $\mathcal{R}_{q}^{2}$)
for $i=0$ to $CH_{in}-1$ do
     for $p=0$ to $K-1$ do
         for $q=0$ to $K-1$ do
              $C_{rot(i,p,q)}$
              $\leftarrow Rot\left(C_{in(i)},I_{in}\times(q+W_{img}\times p)\right)$
         end for
     end for
end for
for $o=0$ to $CH_{out}-1$ do
     $C_{out(o)}$
     $\leftarrow\sum_{i=0}^{CH_{in}-1}\sum_{p=0}^{K-1}\sum_{q=0}^{K-1}Mul\left(C_{rot(i,p,q)},K_{(o,i,p,q)}\right)$
     $C_{out(o)}\leftarrow Add(C_{out(o)},B_{(o)})$
end for
$C_{out}\leftarrow\left(C_{out(0)},\dots,C_{out(CH_{out}-1)}\right)$
$I_{out}\leftarrow I_{in}\times S$
return $C_{out},I_{out}$
The number of multiplications in the convolution layer of UniHENN is $\mathcal{O}(CH_{in}\cdot CH_{out}\cdot K^{2})$, and the number of rotations in the convolution layer of UniHENN is $\mathcal{O}(CH_{in}\cdot K^{2})$. As stated in  [32], the complexity of a single multiplication is $\mathcal{O}(N\cdot D)$, and the complexity of a single rotation is $\mathcal{O}(N\cdot\log N\cdot D^{2})$. Therefore, the complexity of the convolutional layer of UniHENN is $\mathcal{O}(N\cdot D\cdot CH_{in}\cdot K^{2}\cdot(CH_{out}+\log N\cdot D))$.
IV-C Construction of the Average Pooling Layer
CNNs use pooling layers to reduce the input size of convolutional layers. The most commonly used pooling layers are max pooling, min pooling, and average pooling. Max and min pooling require many multiplications [34], while average pooling requires one. Our architecture aims to provide a CNN inference with high speed, we only consider the CNN models without bootstrapping. But, the max pooling and min pooling consume lots of multiplications so it is needed to use the bootstrapping with these pooling layers. Therefore, in UniHENN, we adopt average pooling instead of max pooling and min pooling.
Originally, the average operation is a multiplication of the constant $1/c^{2}$, where $c$ is the size of the kernel. However, in UniHENN, we modify the average pooling layer to eliminate the need for constant multiplication. Our main idea is to apply the constant multiplication in one of the convolutional layers or flatten layers instead of the average pooling layer. It provides the same functionality as the general average pooling but reduces multiplication only changing the order of multiplication of $1/c^{2}$ in the previous layer. So, it can choose lower depth, resulting in efficiency in the overall operation.
The choice is determined by the layer that follows average pooling. If an activation layer follows, the constant multiplication is passed to either the flatten layer or the next convolutional layer, as this layer is linear. Using the linearity $h(x)=Ax+b$, we find $h(x/c^{2})=A(x/c^{2})+b=(1/c^{2})Ax+b$, allowing us to use the weight $(1/c^{2})A$ instead of $A$.
When an activation function such as the square function or the approximate ReLU function follows average pooling, we cannot apply the above logic directly because these functions are not linear. In this case, we can apply the following logic:
  • If the activation function is the square function: Then, $(1/c^{2})^{2}=1/c^{4}$ is applied the following in the convolutional layer or the flatten layer.
  • If the activation function is the approximate ReLU function: Then, we can apply the coefficient of the approximate ReLU ${f(x/c^{2})}=0.375373+(0.5/c^{2})x+(0.117071/c^{4})x^{2}$ if the approximate ReLU is defined as ${f(x)}=0.375373+0.5x+0.117071x^{2}$.
The mechanism of average pooling closely resembles that of the convolutional layer. Suppose an average pooling operation is conducted with a kernel size of $c$. We can then apply a convolutional layer with $c$ as both the kernel size and stride and use a constant multiplication of $1/c^{2}$ as described. The overview of this logic is shown in Figure 5.
However, unlike the convolutional layer, invalid values are introduced in the average pooling layer because the flattened kernels are not multiplied, as illustrated in Figure 6. Due to this issue, the gap between data requires maximum rotation to the left. The exact interval is as follows:
$(W_{img}+1)\times(c-1)$ (2)
IV-D Construction of the Fully Connected (FC) Layer
In an FC layer, the input data is multiplied by a weights matrix, as depicted in Figure 8 (a). PycrCNN performs the same procedure as a plain FC layer by multiplying encrypted data with the weights matrix. However, in this case, matrix multiplication requires significant computational power and time.
(a) PyCrCNN
(b) TenSEAL
Figure 8: Structure of the FC layer of (a) PyCrCNN and (b) TenSEAL.
Employing the vector multiplication technique [18] offers a way to facilitate FC layer operations on the ciphertext. As depicted in Figure 8 (b), TenSEAL employs the diagonal method for vector-matrix multiplication [7, 18]. In this approach, rotated data is multiplied by a rotated weights vector. To utilize vector multiplication, the width of the weight matrix is set to the input size by padding with zeros. Consequently, the number of multiplications and rotations depends on the FC layer’s input size. Generally, the input size of the FC layer is larger than its output size, which may consume unnecessary resources. To minimize the number of resource-intensive operations, including multiplication and rotation, we optimized the logic to depend on the FC layer’s output size.
(a) Step 1
(b) Step 2
(c) Step 3
(d) Step 4
(e) Step 5
Figure 9: Construction of FC layer of UniHENN.
In Figure 9, the overall process of vector multiplication in UniHENN is illustrated. First, as shown in Figure 9 (a), pad zeros until the number of rows is a multiple of the output size, and then rotate the columns of the weight matrix upward, incrementing one column at a time as shown in Figure 9 (b). Next, append the segments of each array with a negative index to the end of the array. Finally, as depicted in Figure 9 (c), divide the matrix lengthwise by the output size and concatenate the parts horizontally; the weight matrix $M_{w}$ then becomes $M_{rot}$ as shown in Figure 9 (d). Note that generating an optimized weight matrix (from Figure 9 (a) to Figure 9 (d)) involves modifying the weights in plaintext. Therefore, the operation time is negligible. Algorithm 2 provides the details of the process for constructing the fully connected layer.
In the inference process, the input ciphertext is rotated and multiplied by the weight matrix $M_{w}$, as depicted in Figure 9 (e). Matrix multiplication can be performed in encrypted form with only multiplication operations that depend on the output size. Generally, since the output size of the FC layer is smaller than the input size, operations can be more efficient than with the diagonal method. For example, if the FC layer has an input size of 64 and an output size of 10, then the total number of multiplications and rotations would be 10, not 64.
Algorithm 2 Construction of the fully connected layer
Input:
- Ciphertext : $C_{in}\in\mathcal{R}_{q}^{2}$
- Weight matrix : $M_{w}\in\mathbb{R}^{DAT_{out}\times DAT_{in}}$
- Bias : $B\in\mathbb{R}^{DAT_{out}}$
Output:
- Ciphertext : $C_{out}\in\mathcal{R}_{q}^{2}$
for $o=0$ to $CH_{out}-1$ do
     $V_{1}\leftarrow M_{rot(o)}[:DAT_{out}-o]+[0]\times o$
     $C_{out(o)}\leftarrow Mul(Rot(C_{in},o),V_{1})$
     $V_{2}\leftarrow[0]\times(DAT_{out}-o)+M_{rot(o)}[DAT_{out}-o:]$
     $V_{3}\leftarrow Mul(Rot(C_{in},o-DAT_{out}),V_{2})$
     $C_{out(o)}\leftarrow Add(C_{out(o)},V_{3})$
end for
for $i=1$ to $\lceil DAT_{in}/DAT_{out}\rceil-1$ do
     $C_{out}\leftarrow Add(C_{out},Rot(C_{out},i\times DAT_{in}))$
end for
$C_{out}\leftarrow Add(C_{out},B)$
return $C_{out}$
The number of multiplications in the fully connected (FC) layer of UniHENN is $\mathcal{O}(DAT_{out})$, and the number of rotations in the FC layer of UniHENN is $\mathcal{O}(DAT_{in}/DAT_{out}+DAT_{out})$. This implies that if $DAT_{in}$ is smaller than the square of $DAT_{out}$, the rotations can be denoted as $\mathcal{O}(DAT_{out})$. In this scenario, the time complexity of the FC layer depends solely on the output size of the FC layer. Thus, the complexity of the FC layer of UniHENN is $\mathcal{O}((DAT_{in}/DAT_{out}+DAT_{out})\cdot N\cdot\log N\cdot D^{2})$.
IV-E Construction of the Flatten Layer
The outputs of the convolutional layer (or average pooling layer) typically have an interval (the reason is detailed in Section IV-B). Additionally, the output of the convolutional layer is usually sparse, as the output from each filter resides in a distinct ciphertext. In a flatten layer, the operation is performed to collect each ciphertext and integrate it into a single ciphertext. Sparse ciphertext leads to unnecessary time and memory consumption. To avoid this, we construct flatten layers to remove invalid data between valid data and to compress all convolutional computation results into a single kernel ciphertext.
IV-E1 Removing the row interval
Figure 10: Removing row interval when the previous layer is either a convolutional layer or an approximate ReLU layer.
After passing through the convolution layer or fully connected layer, the valid data in the output ciphertext will have offset, which we call row interval. We propose an algorithm to remove the row interval for two situations: the preceding layer is an average pooling layer or the preceding layer is not an average pooling. As detailed in Section IV-C, the square layer does not handle constant multiplication, so it is not considered when identifying the previous layer.
As illustrated in Figure 10, if the previous layer is an average pooling layer, the given input ciphertexts have invalid values in the intervals between each set of values. Furthermore, since we deferred the constant multiplication ($1/c^{2}$) in the pooling layer, we need to multiply $1/c^{2}$ to the ciphertexts. Section IV-E2 describes the three steps to address this issue.
If the preceding layer is either a convolutional layer or an approximate ReLU layer, it will contain some zero values between each set of non-zero values. We can exploit this to efficiently execute the flatten operation by stacking the non-zero values using rotation and addition.
Additionally, if $W_{in}=1$, there is only one column. Therefore, this removing row interval step can be skipped.
[…]
IV-E2 Removing invalid data and row interval
Figure 11: Removing invalid data between valid data when the previous layer is an average pooling layer.
[…]
Suppose $S_{total}$ is the value obtained by multiplying the stride values of all convolutional layers by the kernel sizes of all average pooling layers, $W_{img}\times H_{img}$ is the image size and $W_{in}\times H_{in}$ is the output size of the last average pooling layer.
[…]
IV-E3 Removing the column interval
Figure 12: Overall mechanism for compressing all convolutional computation results into a single kernel ciphertext.
[…]
These additional steps help to remove the column intervals effectively, ensuring that the data is adequately flattened and ready for further processing in the CNN pipeline.
IV-E4 Packing multiple ciphertexts
Figure 13: Overall mechanism for multiple ciphertexts into a single ciphertext.
Figure 13 shows the detail of the steps taken to combine the multiple output vectors produced by the convolution operation into a singular ciphertext. The descriptions are as follows:
[…]
After completing this procedure, the outputs are flattened into a single ciphertext. The number of slots occupied in this ciphertext will be $W_{in}\times H_{in}\times CH_{in}$.
The convolutional layer will also require two multiplication operations when the stride $S$ exceeds one. However, if $S_{total}$ equals one, the process illustrated in Figure 12 (a) becomes unnecessary. Thus, only a single multiplication operation would be needed.
The number of multiplications in the flatten layer of UniHENN is $\mathcal{O}(W_{in}+H_{in})$, and the number of rotations in the flatten layer of UniHENN is $\mathcal{O}(CH_{in}\cdot S_{total}+W_{in}+H_{in})$.
Therefore the complexity of flatten layer of UniHENN is $\mathcal{O}((CH_{in}\cdot S_{total}+W_{in}+H_{in})\cdot N\cdot\log N\cdot D^{2})$
IV-F Construction of the Activation Layer
Activation functions require many multiplications [34], so it requires many computing resources to implement them with HE. To address this, we approximate each activation function using a low-degree polynomial. For instance, a 2-degree polynomial approximation of the ReLU function [20] is represented as ${f(x)=0.375373+0.5x+0.117071x^{2}}$. In UniHENN, these approximated activation functions are used instead of the real activation functions for efficiency.
V Optimizing Ciphertext Size for Batch Processing
As shown in Figure 3, UniHENN combines multiple encrypted input data into a single ciphertext for batch processing. However, the ciphertext size for each input should not be allocated solely based on the size of the input data. Instead, it should also consider the intermediate or final output values produced by performing CNN operations. This prevents overlapping issues with other encrypted input data’s intermediate or final output values. We can determine this size in advance by examining the structure of the CNN model.
This section will explain how the output size can be determined in each of the layers (convolutional layer, average pooling layer, flatten layer, and fully connected layer) implemented in UniHENN.
V-A Convolutional Layer
The size of the convolution layer’s output data depends on three parameters: stride, padding, and kernel size.
  • Stride: In the convolutional layer of UniHENN, the interval between each row of data equals the product of the strides from all previous layers, while the interval between each column equals this product multiplied by the image width.
  • Padding: In a convolutional layer with padding, extra space is required to add 0 elements. If there are $L$ convolutional layers with padding spaces $P_{1},P_{2},...,P_{L}$, then allocate a space of length $P=P_{1}+P_{2}+...+P_{L}$ in the input ciphertext. This space is used for each padding process. As padding is accounted for in the initial input image size, it is not considered when calculating the largest size in Theorem 1.
  • Kernel size: A larger kernel size results in a smaller output size, as explained in Theorem 1.
The output size of the convolutional layer does not always exceed the input size when the kernel is greater than or equal to the stride. This is because the convolutional operation is designed to be calculated independently in each channel. The proof of this statement is described in Theorem 1.
In Theorem 1, one condition is that the stride value is not larger than the kernel size. This is realistic, as a larger stride value would result in data loss from the image.
Theorem 1.
Let $W_{img}\times H_{img}$ be the image size and $L$ the number of total convolutional layers. Denote $(W_{in(i)},H_{in(i)})$ and $(W_{out(i)},H_{out(i)})$ as the input and output sizes of the $i$-th convolutional layer, and the kernel size is $(KW_{i},KH_{i})$, stride size is $(SW_{i},SH_{i})$, and padding size is $(PW_{i},PH_{i})$ for all $i\in\left[1,L\right]$. Additionally, let
$SH_{(i)}=\prod\limits_{j=1}^{i}SH_{j}$ (3)
Then, $H_{img}\geq H_{out(i)}\times SH_{(i)}$, where $KH_{i}\geq SH_{i}\geq 1$ for all $i\in\left[1,L\right]$.
Proof.
We obtain that $H_{out(i)}$ :
$H_{out(i)}=\left[\frac{H_{in(i)}-KH_{i}}{SH_{i}}\right]+1$ (4)
Given the hypothesis, the following inequality holds:
$H_{out(i)}\times SH_{i}$ $=\left(\left[\frac{H_{in(i)}-KH_{i}}{SH_{i}}\right]+1\right)\times SH_{i}$ (5)
$=\left[\frac{H_{in(i)}-KH_{i}+SH_{i}}{SH_{i}}\right]\times SH_{i}$ (6)
$\leq\left[\frac{H_{in(i)}}{SH_{i}}\right]\times SH_{i}\leq H_{in(i)}=H_{out(i-1)}$ (7)
When $i>1$, $H_{in(i)}=H_{out(i-1)}$. As proven by mathematical induction, inserting an average pooling layer or an activation function between convolutional layers does not change the size. ∎
Note that after passing through the $i$-th convolutional layer, the interval in each column of data is $W_{img}\times(SH_{(i)}-1)$ (detailed in  IV-B). Given $W_{img}$ columns of data or fewer, the image size of $H_{img}\times W_{img}$ is not exceeded by Theorem 1.
Corollary 1.1.
Denote
$SW_{(i)}=\prod\limits_{j=1}^{i}SW_{j}$ (8)
Then, $W_{img}\geq W_{out(i)}\times SW_{(i)}$, where $KW_{i}\geq SW_{i}\geq 1$ for all $i\in\left[1,L\right]$.
Proof.
The proof is analogous to Theorem 1 and is therefore omitted. ∎
We adopt a slot-based operation approach, allowing data with remaining intervals to pass through to the next layer unchanged. The row interval is defined as the length between row vectors, and the column interval is the length between column vectors. Specifically, if the stride of the $i$-th convolutional layer is $S_{i}$, the interval between each data value after this layer is:
row interval $:\prod\limits_{j=1}^{i}S_{j}-1$ (9)
column interval $:W_{img}\times(\prod\limits_{j=1}^{i}S_{j}-1)$ (10)
By Theorem 1 and Corollary 1.1, each data value stays within the specified index, confirming the flawless operation of the convolutional layer in UniHENN.
V-B Average Pooling Layer
Assuming that the kernel size in the average pooling layer is $c$, each channel uses a filter with a kernel size of $(c,c)$ and a stride of $(c,c)$. The result is then multiplied by $1/c^{2}$. We omit the multiplication of a vector consisting of 0s and 1s, allowing some non-zero but unused values (i.e., invalid values) to overlap with other values. According to Theorem 1, the number of slots in the ciphertext data remains unchanged when passing through UniHENN’s average pooling. However, to prevent invalid values from occupying the positions of other used data, an additional space of $(W_{img}+1)\times(c-1)$ is required (detailed in Section IV-C).
V-C Flatten Layer
Assume that we flatten data from $CH_{in}$ channels, each of size $W_{in}\times H_{in}$. After flattening, the ciphertext data will have $W_{in}\times H_{in}\times CH_{in}$ slots. This size can exceed the maximum size of the layers preceding the flatten layer. We compare these two sizes to determine the largest possible size of the hidden layer. Because the model has plaintext information, comparison operations are possible.
V-D Fully Connected Layer
Let $CH_{in}$ and $CH_{out}$ be the input size and output size of the FC layer, respectively. The size of the slot required in the FC layer is given by:
$CH_{out}\times\left\lceil\frac{CH_{in}}{CH_{out}}\right\rceil$ (11)
Details of how this formula was derived can be found in Section IV-D. The maximum size up to the preceding layer is compared with this value and updated to the larger of the two.
VI Experiments
This section presents experimental results demonstrating the feasibility and effectiveness of UniHENN.
We implement seven distinct CNN models, including LeNet-1 [27] and LeNet-5 [29], to demonstrate that UniHENN can be applicable to a wide range of model architectures.
VI-A Experimental Settings
We used a server service on the NAVER Cloud platform [42] with the following specifications: 16 vCPU cores (Intel(R) Xeon(R) Gold 5220 CPU @ 2.20GHz), 64GB of memory, and a 50GB SSD.
VI-B Datasets
We utilize four different datasets for each model: MNIST [15], CIFAR-10 [26], USPS [19], and ECG [38].
VI-B1 MNIST Dataset
MNIST is a dataset [15] consisting of 70,000 grayscale images of handwritten digits, each with a size of $28\times 28$ pixels. The dataset contains 10 classes, representing the digits from 0 to 9. The dataset consists of two parts: 60,000 data for training and 10,000 data for testing. We use these settings in our experiments.
VI-B2 CIFAR-10 Dataset
CIFAR-10 is a dataset [26] comprising 60,000 RGB images, each with dimensions $3\times 32\times 32$ pixels. Each image is labeled with one of 10 classes. The dataset consists of two parts: 60,000 data for training and 10,000 data for testing. We use these settings in our experiments.
VI-B3 USPS Dataset
The USPS dataset [19] consists of 9,298 grayscale images of handwritten digits, each with dimensions $16\times 16$ pixels. The dataset also contains 10 classes, representing the digits from 0 to 9. The dataset consists of two parts: 7,291 data for training and 2,007 data for testing. We use these settings in our experiments.
VI-B4 ECG Dataset
The ECG dataset from the MIT-BIH arrhythmia database [38] is used for the performance evaluation of our proposed patient-specific ECG approach. It has a total of 109,446 samples with a sampling frequency of 125Hz. The signals represent electrocardiogram (ECG) shapes of heartbeats in normal cases and cases affected by various arrhythmias and myocardial infarction. Each data point is labeled with one of 5 classes. In Abuadbba et al. [3], they pre-processed the ECG dataset to choose a total of 43 out of 48 records are selected and each signal was rescaled by the min-max normalization. And, choose the biorthogonal wavelet as decomposition and reconstruction wavelet function. We use 13,245 data for training and 13,245 data for testing which is in the https://github.com/SharifAbuadbba/split-learning-1D.
We use a preprocessed dataset generated by Abuadbba et al. [3], where each data point is one-dimensional and has 128 columns.
VI-B5 Hyperparameter Settings for Training
We trained all models with the learning rate of 0.001 and the batch size of 32. We trained the MNIST and USPS models for 15 epochs, the CIFAR-10 model for 25 epochs, and the ECG model for 30 epochs.
VI-C Library and Parameter Setup
We used SEAL-Python [2], a library that provides Python APIs for performing homomorphic operations like key generation, encryption, decryption, addition, multiplication, and rotation. The reason for using the SEAL-based library is that baseline models TenSEAL, and PyCrCNN in our experiments, are based on SEAL. So, we use it to balance the experiment. And, we use the pytorch and torchvision library for training. The detailed information of libraries and the experimental setup is described in the GitHub site : https://github.com/hm-choi/uni-henn. The specific parameters used in our experiments are detailed in Table II. All security parameters meet the 128-bit security level as described in [43].
TABLE II: Parameters used in UniHENN. $PK$ and $SK$ represent the size of the public key and the secret key. $GK$ represents the size of the Galois key, $RK$ represents the size of the set of Relinearization keys, and # mult means the total number of allowed multiplications.
# slots 8,192
scale factor 32
log Q 432
$PK$ (MB) 1.87
$SK$ (MB) 0.94
$GK$ (GB) 0.57
$RK$ (MB) 22.52
$ctxt$ (MB) 1.68
# mult (depth) 11
The public key and secret key are also referred to as the encryption key and decryption key, respectively. The set consisting of the Galois key and the relinearization keys is also known as the evaluation key.
VII Experimental Results
In this section, we perform various experiments to verify the feasibility and evaluate the performance of UniHENN. Each experiment in this paper was conducted under identical conditions and repeated 30 times to ensure consistency. The objectives of our experiments are as follows:
  • Comparison of Inference Time with State-of-the-Art Solutions. In this section, we compare the inference performance of UniHENN against two state-of-the-art HE-based deep learning inference frameworks: TenSEAL [7] and PyCrCNN [16]. We selected TenSEAL for comparison because it is a well-known open-source library that uses the im2col algorithm to implement CNNs. This algorithm allows TenSEAL to achieve highly optimized and efficient inference times for convolution operations. However, this efficiency comes at the cost of flexibility: TenSEAL’s architecture only supports models with a single convolutional layer. Therefore, TenSEAL is not a viable option for CNN architectures with multiple convolutional layers. PyCrCNN is an alternative to TenSEAL that does not employ the im2col algorithm, making it applicable to a broader range of CNN architectures. However, this flexibility comes with a significant drawback: PyCrCNN exhibits slower inference times than TenSEAL. Our experiments show that UniHENN strikes a balance between these extremes. It achieves inference times comparable to those of TenSEAL while also supporting CNN architectures with multiple convolutional layers. This makes UniHENN a more versatile solution for HE-based deep learning inference.
  • Adaptability of UniHENN for Various CNN Model Architectures. One of the compelling features of UniHENN is its ability to adapt to a variety of CNN architectures, including both complex and 1D CNN models. To demonstrate this adaptability, we conducted experiments with several CNN models, including LeNet-5, a seminal CNN model widely adopted for digit recognition tasks [51, 36]. LeNet-5 is more complex than many other models, with multiple convolutional layers and fully connected layers. Our successful implementation of LeNet-5 in UniHENN highlights the system’s ability to handle large and complex CNN models effectively. We also consider a 1D CNN architecture particularly useful for sequence-based tasks such as time-series analysis, natural language processing, and certain bioinformatics applications. This capability sets UniHENN apart from solutions like TenSEAL, which is constrained to supporting only single convolutional layer models and does not offer support for 1D CNNs. Through these experiments, we aim to demonstrate UniHENN’s comprehensive applicability and its ability to adapt to various CNN architectures, making it a versatile tool for secure and efficient deep learning inference across diverse application domains.
  • Adaptability of UniHENN with Various Datasets. One of the key strengths of UniHENN is its adaptability to various kinds of data. To demonstrate this, we conducted experiments using four diverse datasets: MNIST, CIFAR-10, USPS, and ECG. MNIST and CIFAR-10 are widely used image classification datasets that serve as standard benchmarks in the deep learning community. MNIST consists of grayscale images of handwritten digits, while CIFAR-10 comprises coloured images spanning ten different object classes. USPS is another grayscale image dataset used for handwriting recognition. ECG represents electrocardiogram data and is commonly used in healthcare applications for diagnosing various heart conditions. By demonstrating that UniHENN performs well across these varied datasets, we aim to show its wide applicability to tasks involving images of different sizes and complexities, as well as specialized domains like healthcare.
VII-A Experiment 1: Comparison of Inference Time between UniHENN, TenSEAL, and PyCrCNN
In this experiment, we implement the model, denoted as $M_{1}$, using UniHENN, TenSEAL, and PyCrCNN. The $M_{1}$ model was introduced in the TenSEAL paper [7] and is well-suited for implementation using TenSEAL. Despite its simple architecture, which contains only one convolutional layer, the model achieves a high accuracy of 97.65%. Table III provides detailed specifications for this model architecture. The Conv2D parameters–$ch_{in}$, $ch_{out}$, size, and stride–represent the number of input channels, the number of output channels, kernel size, and stride for a two-dimensional convolutional layer, respectively. Similarly, $data_{in}$ and $data_{out}$ in the FC layer indicate the input and output dimensions. We use these notations to represent all other remaining models presented in this paper.
TABLE III: Detailed parameters for $M_{1}$.
Layer Parameter # Mult
Conv2d $ch_{in}$=1, $ch_{out}$=8, size=4, stride=3 1
Square - 1
Flatten - 2
FC1 $data_{in}$ = 648, $data_{out}$ = 64 1
Square - 1
FC2 $data_{in}$ = 64, $data_{out}$ = 10 1
Total #Mult - 7
Table IV presents the results. In terms of inference time for a single sample, the TenSEAL implementation exhibited the highest performance, taking 6.298 seconds. TenSEAL achieved this efficiency by optimizing convolution operations on the input data using the im2col algorithm. UniHENN took 16.248 seconds, making it 2.58 times slower than TenSEAL. PyCrCNN was the slowest, requiring 154.494 seconds. However, UniHENN has the advantage when handling multiple data points simultaneously. It outperforms TenSEAL by supporting batch operations for up to ten input samples, allowing concurrent inference calculations. TenSEAL and PyCrCNN do not offer parallel processing for multiple samples, resulting in total times that increase proportionally with the number of input samples. To process ten input samples together, UniHENN still takes 16.248 seconds, approximately 3.9 times and 94.6 times faster than TenSEAL’s 63.706 seconds and PyCrCNN’s 1537.228 seconds, respectively.
TABLE IV: Comparison of the average (with standard deviation) inference time in seconds on the MNIST dataset between PyCrCNN and UniHENN for the $M_{1}$ model architecture.
Layer TenSEAL PyCrCNN UniHENN
# of samples 1 10 1 10 1 & 10
Drop DEPTH - - - -
0.065
Conv2d
2.522
25.202
31.844
316.536
3.436
Square
0.024
0.242
25.082
247.291
0.281
Flatten - -
0.000
0.000
5.285
FC1
3.540
36.147
94.883
946.841
6.783
Square
0.014
0.140
1.527
15.017
0.011
FC2
0.198
1.975
1.158
11.542
0.386
Total
6.298
63.706
154.494
1537.228
16.248
We conducted additional experiments to determine the number of samples at which UniHENN starts to outperform TenSEAL. The experiment was carried out by incrementally increasing the number of samples and observing the inference time. The results are presented in Figure 14.
Figure 14: Average inference time (in seconds) for the $M_{1}$ model architecture using UniHENN and TenSEAL with varying the number of input samples.
In Figure 14, the inference times for UniHENN and TenSEAL using the $M_{1}$ model are presented. Since TenSEAL does not support batched inference, its inference time increases linearly as the number of input images grows. In contrast, UniHENN does support batched inference, allowing up to 10 MNIST images to be processed simultaneously in the $M_{1}$ model. The figure reveals that UniHENN’s inference time becomes shorter than TenSEAL’s starting at a batch size of 3. This demonstrates that UniHENN surpasses TenSEAL in efficiency when concurrently processing $k$ images, where $k\geq 3$.
VII-B Experiment 2: Comparison of Inference Time between UniHENN and PyCrCNN for LeNet-1
In this experiment, we utilize the LeNet-1 model, denoted as $M_{2}$. The model accuracy of $M_{2}$ is 98.62%, slightly higher than $M_{1}$. The detailed specifications of these model architectures can be found in Table V. Implementing the hyperbolic tangent function (tanh) in HE is computationally challenging; therefore, we have substituted the activation function from the tanh to the square function. We find that the modified LeNet-1 model with the square activation achieves an accuracy of 98.62% comparable to the 98.41% accuracy of the original LeNet-1 model when tested on a 10,000 sample MNIST dataset. This indicates that the modification in the activation function does not significantly impact the model’s accuracy.
TABLE V: Detailed parameters for $M_{2}$.
Layer Parameter #Mult
Conv2d $ch_{in}$=1, $ch_{out}$=4, size=5, stride=1 1
Square - 1
AvgPool2d kernel size=2 0
Conv2d $ch_{in}$=4, $ch_{out}$=12, size=5, stride=1 1
Square - 1
AvgPool2d kernel size=2 0
Flatten - 2
FC1 $data_{in}$=192, $data_{out}$=10 1
Total #Mult - 7
This experiment aims to confirm that UniHENN can support CNN models with more than two convolutional layers. Importantly, the TenSEAL library cannot be used in this experiment due to its constraint of supporting only a single convolutional layer, a limitation arising from its im2col algorithm implementation. The results are presented in Table VI.
TABLE VI: Comparison of the average (with standard deviation) inference time in seconds on the MNIST dataset between PyCrCNN and UniHENN for $M_{2}$ model architecture.
Model PyCrCNN UniHENN
Drop DEPTH - 0.066 (0.003)
Conv2d 169.182 (0.891) 3.715 (0.048)
Square 88.759 (0.605) 0.140 (0.003)
AvgPool2d 1.955 (0.080) 0.535 (0.012)
Conv2d 520.452 (1.609) 21.697 (0.150)
Square 13.297 (0.116) 0.260 (0.005)
AvgPool2d 0.419 (0.019) 0.875 (0.011)
Flatten 0.000 (0.000) 2.279 (0.020)
FC1 6.527 (0.070) 0.522 (0.014)
Total 794.064 (2.279) 30.090 (0.155)
In Table VI, UniHENN takes 30.090 seconds, making it approximately 26.4 times faster than PyCrCNN, which takes 794.064 seconds. In this experiment, most of the computational time for both UniHENN and PyCrCNN is consumed in the second convolutional layer. These findings highlight the importance of optimizing convolutional layer operations for time-efficient CNN inference in the context of HE.
Furthermore, both Experiments 1 and 2 employ the same input ciphertext, showcasing that UniHENN enables diverse CNN models without requiring re-encryption, provided the supported HE parameters across the models are identical.
VII-C Experiment 3: Adaptibility of UniHENN for CNN Models with Approximate ReLU activation
In this experiment, we implement a CNN model, denoted as $M_{3}$, with approximate ReLU for the MNIST dataset. Originally, we planned to modify $M_{2}$ by replacing the square activation function with approximate ReLU, but the performance was not satisfactory. Therefore, we redesigned the model to achieve better performance.
The model accuracy of $M_{3}$ is 98.22%, which is similar to $M_{2}$’s accuracy of 98.62%. Table VII provides detailed specifications for this model architecture.
TABLE VII: Detailed parameters for $M_{3}$.
Layer Parameter #Mult
Conv2d $ch_{in}$=1, $ch_{out}$=6, size=3, stride=1 1
Approx ReLU $f(x)=0.375373+0.5x+0.117071x^{2}$ 2
AvgPool2d kernel size=2 0
Flatten - 2
FC1 $data_{in}$ = 1014, $data_{out}$ = 120 1
Approx ReLU $f(x)=0.375373+0.5x+0.117071x^{2}$ 2
FC2 $data_{in}$ = 120, $data_{out}$ = 10 1
Total #Mult - 9
Table VIII shows a layer-by-layer breakdown of the inference time for the $M_{3}$ model architecture. The Flatten and FC1 layers are the most time-consuming, taking an average of 9.603 seconds and 15.557 seconds, respectively. These two layers alone contribute significantly to the total inference time of 29.104 seconds. While UniHENN is significantly slower than non-HE models, it provides enhanced privacy and security, which may be crucial for specific applications or compliance requirements. The inference time of 29.104 seconds is still under 30 seconds, suggesting that UniHENN is practical for real-world applications, especially in contexts where data security is paramount. This inference time could be considered acceptable depending on the specific use case and the sensitivity of the data being processed. These results further validate the adaptability and efficiency of UniHENN in handling various CNN architectures with customized functionalities like approximate ReLU.
TABLE VIII: Average (with standard deviation) inference time in seconds on the MNIST dataset for the $M_{3}$ model architecture.
Model UniHENN
Drop DEPTH 0.035 (0.001)
Conv2d 1.863 (0.029)
Approx ReLU 0.513 (0.011)
AvgPool2d 1.030 (0.024)
Flatten 9.603 (0.081)
FC1 15.557 (0.111)
Approx ReLU 0.045 (0.002)
FC2 0.459 (0.010)
Total 29.104 (0.171)
VII-D Experiment 4: Adaptibility of UniHENN for LeNet-5
In this experiment, we implement the LeNet-5 model [29], denoted as $M_{4}$, with the only modification being the activation function, which is changed from tanh to square. The model accuracy of $M_{4}$ is 98.91%. Table IX provides detailed specifications for this model architecture.
TABLE IX: Detailed parameters for $M_{4}$.
Layer Parameter #Mult
Conv2d $ch_{in}$=1, $ch_{out}$=6, size=5, stride=1 1
Square - 1
AvgPool2d kernel size=2 0
Conv2d $ch_{in}$=6, $ch_{out}$=16, size=5, stride=1 1
Square - 1
AvgPool2d kernel size=2 0
Conv2d $ch_{in}$=16, $ch_{out}$=120, size=5, stride=1 1
Square - 1
Flatten - 0
FC1 $data_{in}=120$, $data_{out}=84$ 1
Square - 1
FC2 $data_{in}=84$, $data_{out}=10$ 1
Total #Mult - 9
Table X shows a layer-by-layer breakdown of the inference time for the $M_{4}$ model architecture. The results show that the inference time for $M_{4}$ is 740.127 seconds. This is substantially slower than the previous experiments, which could be attributed to the model’s increased complexity. The slow performance emphasizes the need for further optimization, especially if UniHENN is to be broadly applied to more complex CNN architectures like LeNet-5 for real-world applications. Note that the convolutional layers are the most time-consuming, highlighting the critical need for optimizing these operations when implementing CNNs using HE. With such a long inference time, the immediate practicality of using this model for real-time or near-real-time applications is limited. However, this could be acceptable for services that require strong privacy controls and where data security is a higher priority than speed. For example, in healthcare or financial services, where data may be extremely sensitive, this level of privacy may justify the slower inference times.
TABLE X: Average (with standard deviation) inference time in seconds on the MNIST dataset for the $M_{4}$ model architecture.
Layer UniHENN
Drop DEPTH 0.035 (0.002)
Conv2d 5.244 (0.048)
Square 0.312 (0.005)
AvgPool2d 0.865 (0.017)
Conv2d 48.222 (0.266)
Square 0.567 (0.010)
AvgPool2d 1.453 (0.020)
Conv2d 668.688 (1.241)
Square 2.613 (0.030)
Flatten 3.995 (0.044)
FC1 7.669 (0.074)
Square 0.011 (0.000)
FC2 0.454 (0.009)
Total 740.127 (1.381)
VII-E Experiment 5: Adaptability of UniHENN for CNN Models on Color Images
In this experiment, we evaluate the adaptability of UniHENN using the CIFAR-10 color image dataset. We implement a CNN model, denoted as $M_{5}$, which is a modified version of $M_{4}$, to achieve satisfactory accuracy on CIFAR-10. The model accuracy of $M_{5}$ is 73.26%. Table XI provides detailed specifications for this model architecture.
Note that a comparison with TenSEAL is not feasible, as TenSEAL does not process multiple channels, which is essential to process color images. Additionally, we attempted an experiment with PyCrCNN under the same settings but failed to obtain results despite running the experiment for approximately 15 hours. This failure is attributed to PyCrCNN’s approach of encrypting each parameter with an individual ciphertext, which demands substantial memory and computational time.
TABLE XI: Detailed parameters for $M_{5}$.
Layer Parameter #Mult
Conv2d $ch_{in}$=3, $ch_{out}$=16, size=3, stride=1 1
Square - 1
AvgPool2d kernel size=2 0
Conv2d $ch_{in}$=16, $ch_{out}$=64, size=4, stride=1 1
Square - 1
AvgPool2d kernel size=2 0
Conv2d $ch_{in}$=64, $ch_{out}$=128, size=3, stride=1 1
Square - 1
AvgPool2d kernel size=4 0
Flatten - 1
FC1 $data_{in}=128$, $data_{out}=10$ 1
Total #Mult - 8
Table XII presents the experimental results, showing that the total inference time for UniHENN is approximately 20 minutes. While this may seem long, it is important to note that the convolutional layers are particularly time-consuming, requiring about 256.000 seconds for the first layer and 938.446 seconds for the second. This is consistent with the results of previous experiments. Despite the long inference time, we consider it tolerable given the inherent complexities of performing inference on color images, a capability not offered by alternative solutions.
TABLE XII: Average (with standard deviation) inference time in seconds for the $M_{5}$ model architecture.
Layer UniHENN
Drop DEPTH 0.155 (0.004)
Conv2d 8.935 (0.062)
Square 0.703 (0.011)
AvgPool2d 1.891 (0.022)
Conv2d 256.000 (0.867)
Square 1.839 (0.023)
AvgPool2d 4.466 (0.049)
Conv2d 938.446 (1.884)
Square 2.064 (0.019)
AvgPool2d 4.466 (0.049)
Flatten 33.370 (0.162)
FC1 3.146 (0.025)
Total 1251.444 (2.531)
VII-F Experiment 6: Comparison of Inference Time of UniHENN on Grayscale Images
To assess the adaptability of UniHENN to diverse datasets, we performed experiments using the USPS dataset and the $M_{6}$ model architecture. For performance comparison, we also implemented the same model using PyCrCNN. The model accuracy of $M_{6}$ is 94.21%. Table XIII provides detailed specifications for this model architecture.
TABLE XIII: Detailed parameters for $M_{6}$.
Layer Parameter #Mult
Conv2d $ch_{in}$=1, $ch_{out}$=6, size=4, stride=2 1
Square - 1
Flatten - 2
FC1 $data_{in}$ = 294, $data_{out}$ = 64 1
Square - 1
FC2 $data_{in}$ = 64, $data_{out}$ = 10 1
Total #Mult - 7
Table XIV shows that UniHENN is approximately 6 times faster than PyCrCNN, with a total inference time of 12.307 seconds compared to PyCrCNN’s 71.139 seconds. Interestingly, the FC1 layer in PyCrCNN consumes a significant portion of the time (42.792 seconds), likely due to its large input size of 294. In contrast, although UniHENN also incurs a relatively high computational cost at the FC1 layer, it is substantially less than that of PyCrCNN. This suggests that UniHENN can efficiently handle FC layers with large input sizes.
TABLE XIV: Comparison of the average (with standard deviation) inference time in seconds on the USPS dataset between PyCrCNN and UniHENN for the $M_{6}$ model architecture.
Layer PyCrCNN UniHENN
Drop DEPTH - 0.066 (0.002)
Conv2d 14.441 (0.125) 2.662 (0.029)
Square 11.238 (0.090) 0.212 (0.004)
Flatten 0.000 (0.000) 3.026 (0.032)
FC1 42.792 (0.121) 5.948 (0.074)
Square 1.502 (0.026) 0.011 (0.000)
FC2 1.166 (0.015) 0.384 (0.009)
Total 71.139 (0.198) 12.307 (0.092)
VII-G Experiment 7: Comparison of Inference Time for a 1D CNN Model between UniHENN and PyCrCNN
To evaluate the adaptability of UniHENN for 1D CNN models, we implemented a 1D CNN model, denoted as $M_{7}$, for processing ECG data. This model, a modified version of the 1D CNN model by Abuadbba et al. [3], achieves an accuracy of 96.87%, which is comparable to the 98.90% achieved by Abuadbba et al.’s original model. For performance comparison, the same model was also implemented using PyCrCNN. Table XV provides detailed specifications for this model architecture.
TABLE XV: Detailed parameters for $M_{7}$.
Layer Parameter #Mult
Conv1d $ch_{in}$=1, $ch_{out}$=2, size=2, stride=2 1
Square - 1
Conv1d $ch_{in}$=2, $ch_{out}$=4, size=2, stride=2 1
Flatten - 1
FC1 $data_{in}$ = 128, $data_{out}$ = 32 1
Square - 1
FC2 $data_{in}$ = 32, $data_{out}$ = 5 1
Total #Mult - 7
Table XVI shows that UniHENN is approximately 3.03 times faster than PyCrCNN, recording a total inference time of 5.119 seconds compared to PyCrCNN’s 15.513 seconds. This indicates that UniHENN is also more efficient even for relatively smaller models, such as 1D CNNs, where the computational time for convolutional layers is not as extensive. This efficiency positions UniHENN as an ideal choice for privacy-sensitive applications like disease diagnosis systems, aligning well with the principles of HE.
TABLE XVI: Average (with standard deviation) inference time in seconds for the $M_{7}$ model architecture on the ECG dataset.
Layer PyCrCNN UniHENN
Drop DEPTH - 0.069 (0.007)
Conv1d 0.556 (0.011) 0.113 (0.003)
Square 4.913 (0.043) 0.075 (0.002)
Conv1d 1.081 (0.041) 0.283 (0.006)
Flatten 0.000 (0.000) 1.622 (0.015)
FC1 8.163 (0.067) 2.765 (0.022)
Square 0.559 (0.011) 0.012 (0.000)
FC2 0.242 (0.006) 0.180 (0.016)
Total 15.513 (0.114) 5.119 (0.039)
VII-H Validation of inference results on encrypted data
The CKKS scheme operates on approximate complex arithmetic, which can introduce minor errors after homomorphic operations. Therefore, it is crucial to validate UniHENN’s inference results by comparing them with the outcomes of the original plaintext inference.
We selected $M_{4}$ and $M_{5}$ as representative models for validation, as the errors were minimal in other models. We evaluated the $M_{4}$ model with 2,000 samples from the MNIST dataset and the $M_{5}$ model with 2,000 samples from the CIFAR-10 dataset. The results showed that all outputs produced by UniHENN were equivalent to those of the original models without any significant loss in accuracy. This indicates that UniHENN can perform secure and highly accurate inferences under the parameter configurations presented in Table II.
VII-I Impact of HE Parameters
We conducted an additional analysis to investigate how the inference time and encrypted inference error vary depending on the parameters of the CKKS scheme, the HE algorithm we employ.
The CKKS scheme’s operation time and decrypted result accuracy are influenced by several parameters, specifically # slots, scale, and depth. Therefore, the inference time for CNN models is intricately linked to these parameters. A detailed analysis of these parameter settings is essential to optimize inference time while maintaining result accuracy.
The # slots parameter is determined by the degree $N$ of the polynomial ring, where $N$ is half the ring’s value. The scale parameter represents the precision of floating-point digits, while depth is influenced by both scale and the polynomial $P$.
VII-I1 Influence of Depth on Inference Time
The operation time increases as the depth increases while using a fixed scale. Figure 15 shows that, in the $M_{1}$ model, the inference time increases linearly with depth. This is attributed to the ciphertext size being dependent on depth, thereby increasing the computational workload. To achieve optimal performance, the depth should align with the number of required multiplications for the model. However, in our architecture, depth is predetermined before model selection. To address this, we set the depth as high as possible and then optimize by fine-tuning the input ciphertext.
Figure 15: Average inference time (seconds) for $M_{1}$ with depth.
VII-I2 Impact of Scale on Error
We measured the error in the $M_{1}$ model’s results as a function of scale. Detailed findings are presented in Figure 16. Increasing scale logarithmically reduces the error. Although a higher scale is advantageous, it consequently leads to a lower depth, given their inverse relationship due to the fixed log Q parameter in the CKKS scheme. Therefore, selecting an optimal scale is crucial for minimizing error while ensuring sufficient depth for CNN inference. A scale of 32 ensures 32-bit decimal point precision. From our observations, we note that the error converges toward zero as the scale increases. Specifically, using a scale value of 30 or higher can significantly reduce the error.
Figure 16: Error in $M_{1}$ with scale.
VIII Limitations
We propose an optimized CNN model inference mechanism based on HE, UniHENN. UniHENN uses batching to reduce inference time, but it has three limitations.
As deep learning has evolved, it has been used in many applications beyond image classification [48], such as detecting attacks on the web [45] or analyzing resource extraction [47]. These DNN models require a large number of operations, especially multiplication operations. To apply homomorphic encryption, multiplication operations must be used within a certain number of times depending on the parameter values used, or bootstrapping must be used. Due to the nature of deep learning models with a large number of multiplication operations, bootstrapping is required, but bootstrapping has the limitation of slowing down the computation speed significantly.
UniHENN does not support multi-core architecture or GPUs. As we tried to deploy our source codes, we chose to use SEAL-Python, an open-source homomorphic encryption framework. But most open-source homomorphic encryption frameworks, including SEAL-Python, are suitable for only CPU and single-thread architecture. Nowadays, some studies for optimizing HE operations in the GPU environment [6, 40] are conducted. The main idea of the study is to improve performance of the HE-operation using the parallelism design. So, to address these limitations, we plan to extend our architecture that support the multi-core environment and GPU environment in the software side with the previous works for future work.
Since UniHENN utilizes the batch system to reduce computational time. UniHENN will be efficient when analyzing sensitive data at scale or when clients infer multiple data at once. However, UniHENN may not be efficient when it is needed to infer single data because the design point of UniHENN is focused on batch operations. To overcome this limitation, we plan to improve the computational efficiency that the operations constituting the convolutional layer or FC layer operation will be distributed in the space where the batch operation was performed to increase efficiency.
IX Related Work
Dowlin et al. [17] proposed CryptoNets, the first HE-based CNN inference scheme. It utilizes square activation instead of ReLU to reduce the total computation cost with a low number of multiplications. However, the input ciphertext size and weights of CryptoNets are very large because they encrypt each pixel of the input image and each element of the weights individually. Chillotti et al. [12] introduced an enhanced version of CryptoNets, applying low-degree polynomials for functions like ReLU, Softplus, and Swish and incorporating techniques like pruning and quantization for more efficient homomorphic operations. Ishiyama et al. [20] adopted an approximation of Google’s Swish activation function using two and four-degree polynomials, coupled with batch normalization, aiming to minimize the accuracy gap between traditional and polynomial-based activation functions. In 2021, Benaissa et al. presented TenSEAL [7], a prominent open-source tool that supports HE-based CNN inference. It utilizes the im2col method to efficiently compute the convolutional layer. However, it cannot support multiple convolutional layers and does not support batch operation. In this paper, we propose UniHENN to address these issues.
Aharoni et al. [4] presented an encryption data structure, the tile tensor, for efficiently calculating deep neural networks containing CNNs. It splits and encrypts the input data considering the model structure. However, the encryption of input data in this scheme depends on the model architecture; sharing the client’s ciphertext with a third party, such as a public cloud service, is challenging because the CNN model’s structure is usually undisclosed. Lee et al. [32] proposed a novel low-latency model optimization strategy using an enhanced HerPN block and masking convolution, significantly reducing multiplication operations in the convolution layer. However, their approach does not consider the design of FC layers. In this paper, we present how to implement all necessary layers for CNNs in HE efficiently.
X Conclusion
In this paper, we introduce a privacy-preserving machine learning framework, UniHENN, addressing a significant challenge in cloud computing. UniHENN enables the construction of HE-enabled CNN models without relying on the traditional im2col operation, enhancing compatibility with diverse ML models and reinforcing privacy in cloud-based machine learning.
Our extensive evaluation of UniHENN across four public datasets using seven different CNNs, including six 2D CNNs and a 1D CNN, demonstrates its robustness and efficacy. Notably, UniHENN maintains accuracy levels comparable to unencrypted models, highlighting its practical applicability without performance compromise.
A key feature of UniHENN is its batch mechanism support, significantly accelerating processing times for multiple datasets concurrently, a breakthrough for cloud computing. Our findings indicate that a CNN constructed with UniHENN outperforms the model $M1$ (comprising one convolutional layer and two fully connected layers) by almost 3.921 times using the TenSEAL library, demonstrating a substantial improvement in processing speed, crucial for real-world applications requiring rapid data processing. We publicly release UniHENN’s source code at https://github.com/hm-choi/uni-henn to enhance its impact, promoting widespread use.
In the future, we aim to optimize UniHENN for multi-thread architectures, leveraging the additional space available in batch computations. This optimization will involve dividing the operations of each layer to enable simultaneous processing, thereby enhancing the efficiency of UniHENN. We aim to establish UniHENN as a practical tool for secure and efficient cloud computing in machine learning applications.

### v3

@@
-UniHENN: Designing More Versatile Homomorphic Encryption-based CNNs without im2col
+
 Abstract
-Homomorphic encryption enables computations on encrypted data without decryption, which is crucial for privacy-preserving cloud services.
-However, deploying convolutional neural networks (CNNs) with homomorphic encryption encounters significant challenges, particularly in converting input data into a two-dimensional matrix for convolution, typically achieved using the im2col technique.
-While efficient, this method limits the variety of deployable CNN models due to compatibility constraints with the encrypted data structure.
-UniHENN, a homomorphic encryption-based CNN architecture, eliminates the need for im2col, ensuring compatibility with a diverse range of CNN models using homomorphic encryption.
-Our experiments demonstrate that UniHENN surpasses the leading 2D CNN inference architecture, PyCrCNN, in inference time, as evidenced by its performance on the LeNet-1 dataset, where it averages 30.090 seconds–significantly faster than PyCrCNN’s 794.064 seconds.
-Furthermore, UniHENN outperforms TenSEAL, which employs im2col, in processing concurrent images, an essential feature for high-demand cloud applications.
-The versatility of UniHENN is proven across various CNN architectures, including 1D and six different 2D CNNs, highlighting its flexibility and efficiency.
-These qualities establish UniHENN as a promising solution for privacy-preserving, cloud-based CNN services, addressing the increasing demand for scalable, secure, and efficient deep learning in cloud computing environments.
+Homomorphic encryption (HE) enables privacy-preserving deep learning by allowing computations on encrypted data without decryption.
+However, deploying convolutional neural networks (CNNs) with HE is challenging due to the need to convert input data into a two-dimensional matrix for convolution using the im2col technique, which rearranges the input for efficient computation.
+This restricts the types of CNN models that can be used since the encrypted data structure must be compatible with the specific model.
+UniHENN is a novel HE-based CNN architecture that eliminates the need for im2col, enhancing its versatility and compatibility with a broader range of CNN models.
+UniHENN flattens input data to one dimension without using im2col.
+The kernel performs convolutions by traversing the image, using incremental rotations and structured multiplication on the flattened input, with results spaced by the stride interval.
+Experimental results show that UniHENN significantly outperforms the state-of-the-art 2D CNN inference architecture named PyCrCNN in terms of inference time.
+For example, on the LeNet-1 model, UniHENN achieves an average inference time of 30.089 seconds, about 26.6 times faster than PyCrCNN’s 800.591 seconds.
+Furthermore, UniHENN outperforms TenSEAL, an im2col-optimized CNN model, in concurrent image processing.
+For ten samples, UniHENN (16.247 seconds) was about 3.9 times faster than TenSEAL (63.706 seconds), owing to its support for batch processing of up to 10 samples.
+We demonstrate UniHENN’s adaptability to various CNN architectures, including a 1D CNN and six 2D CNNs, highlighting its flexibility and efficiency for privacy-preserving cloud-based CNN services.
 Index Terms:
-homomorphic encryption, privacy-preserving machine learning, convolutional neural networks, data privacy
+homomorphic encryption, privacy-preserving machine learning, data privacy
+[…]
+=-15pt
 I Introduction
-The growing use of machine learning (ML) in diverse domains such as finance, autonomous driving, industrial automation, computer vision, speech recognition, and healthcare has raised concerns regarding data privacy and security [35, 50].
-These concerns are particularly relevant for cloud-based ML services, where sensitive data is processed on remote servers [3].
-Homomorphic encryption (HE) is a form of encryption that allows some computations on ciphertext without decrypting it.
-Therefore, it can be used as a powerful tool for preserving privacy in ML applications [22, 23, 20, 33, 13, 24, 30, 31, 11], enabling the confidentiality of sensitive information even when processed by a third party.
-HE is particularly relevant for cloud-based ML services, where sensitive data is often transmitted to remote servers for processing.
-With HE, cloud providers can ensure data privacy in their possession.
-For instance, a bank could employ HE to build an ML model for detecting fraudulent transactions without accessing customer data.
-Similarly, a healthcare provider could analyze patients’ medical records using HE without disclosing the patient’s identity.
-This potential has attracted interest from cloud service providers (e.g., Amazon AWS, Microsoft Azure, Naver Cloud) in integrating HE into ML-based cloud services, such as Data Box Frame [41].
-While convolution operations are fundamental to convolutional neural networks (CNNs), they demand substantial computational resources when performed on ciphertexts within an HE framework.
+The widespread adoption of deep learning has expanded beyond traditional tasks like image classification [50] to diverse applications such as attack detection [47] and resource extraction analysis [49], with machine learning (ML) now playing a crucial role in sensitive sectors including finance, healthcare, and autonomous systems.
+However, this proliferation has raised significant privacy concerns [36, 53], particularly for cloud-based ML services processing sensitive data on remote servers [4].
+Balancing ML’s powerful capabilities with the need to safeguard individual privacy has become a critical challenge, highlighting the urgent need for advanced privacy-preserving techniques in ML applications handling sensitive information.
+Homomorphic encryption (HE) is a powerful tool for preserving privacy in ML applications [25, 22, 34, 14, 27, 26].
+For example, Choi et al.
+introduced privacy-preserving biometric authentication systems [12, 11] that leverage HE for web and cloud environments.
+HE allows computations to be performed on encrypted data without decryption, ensuring that sensitive information remains confidential even when processed by a third party.
+This is particularly beneficial for cloud-based ML services, where sensitive data is often transmitted to remote servers for processing.
+By using HE, cloud providers can ensure data privacy, such as a bank training an ML model to detect fraudulent transactions without accessing customer data or a healthcare provider analyzing patient medical records without disclosing patient identities.
+Additionally, faster inference time with privacy preservation is crucial for real-time applications in these sensitive fields, where quick and secure data processing can significantly improve decision-making and user experience.
+Convolution operations are essential to Convolutional Neural Networks (CNNs) but require significant computational resources when performed on ciphertexts in HE.
 Each kernel shift requires element-wise multiplications and additions, resource-intensive operations on ciphertexts.
@@
 Therefore, encrypting data without restricting it to a specific model would offer greater flexibility and utility.
-In this paper, we introduce UniHENN, a privacy-preserving CNN model built using HE.
-UniHENN enables efficient CNN inference independent of the im2col function, resulting in encrypted input data that is usable across other ML models.
-We devise an effective computation method for convolution operations on flattened input within UniHENN, offering flexible reuse across various ML models.
-Furthermore, UniHENN supports batch operations, thereby facilitating the efficient processing of multiple ciphertexts.
-Our key contributions are summarized as follows:
-  • We introduce UniHENN, a novel and efficient CNN model inference mechanism based on HE, which facilitates input ciphertext reusability.
+We introduce UniHENN, a privacy-preserving CNN model using HE that enables efficient CNN inference without relying on the im2col function.
+Our novel convolution algorithm flattens input data into a one-dimensional form, eliminating the need for data rearrangement.
+The kernel traverses the input, performing convolutions with incremental rotations and structured multiplication, ensuring only relevant elements are used.
+This approach overcomes im2col limitations by reducing image size dependency and focusing on kernel size, improving efficiency and flexibility across various ML models.
+UniHENN also supports batch operations for efficient multi-ciphertext processing.
+Our key contributions include:
+  • We introduce UniHENN, a novel CNN model inference mechanism based on HE, which facilitates input ciphertext reusability.
 Unlike other CNN implementations that require a specific model input structure for the im2col function, UniHENN is designed to handle model-free input ciphertexts, enabling the use of encrypted input data across various HE-based ML services without the need for re-encryption.
@@
 The source code for UniHENN is available at https://github.com/hm-choi/uni-henn.
-  • We empirically demonstrate that UniHENN is efficient.
-Experimental results indicate an average inference time of 30.090 seconds, significantly outperforming the state-of-the-art HE-based 2D CNN inference architecture PyCrCNN [16], which requires an average of 794.064 seconds for inference.
+  • We empirically demonstrate the efficiency of UniHENN.
+Experimental results indicate an average inference time of 30.089 seconds, significantly outperforming the state-of-the-art HE-based 2D CNN inference architecture PyCrCNN [17], which requires an average of 800.591 seconds for inference.
   • We introduce a batch processing strategy for UniHENN to handle multiple data instances in a single operation.
-This strategy efficiently combines multiple data instances into a single ciphertext, reducing the inference time for 10 MNIST images to 16.248 seconds.
+This strategy efficiently combines multiple data instances into a single ciphertext, reducing the inference time for 10 MNIST images to 16.247 seconds.
 It outperforms TenSEAL’s CNN model [7], which takes 63.706 seconds.
-While UniHENN is less efficient than TenSEAL for processing a single image, it surpasses TenSEAL’s performance when processing $k$ images simultaneously, where $k\geq 3$.
+While UniHENN is less efficient than TenSEAL for processing a single image, it surpasses TenSEAL’s performance when processing multiple images simultaneously, particularly when $k\geq 3$.
 II Background
@@
 Convolutional Neural Networks (CNNs) were first introduced by LeCun et al.
-[28] for processing grid-structured data, such as images.
-The LeNet-1 model, presented in 1989 [27], comprised two convolutional layers, two average pooling layers, and one fully connected layer, marking the first concept of the LeNet architecture.
-In 1998, the LeNet-5 model was introduced [29], consisting of three convolutional layers, two pooling layers, and two fully connected layers.
-It was designed to recognize handwritten postal codes efficiently, outperforming traditional methods like the multilayer perceptron model.
-CNNs gained significant traction in the 2010s, primarily due to their exceptional performance on large datasets like ImageNet [14], which includes millions of images across 1,000 classes.
-Currently, research into the utilization of CNNs, such as DGCNN and its applications, is ongoing [47].
-A typical CNN includes an input layer, multiple hidden layers, and an output layer.
+[31] for processing grid-structured data like images.
+In 1989, the LeNet-1 model was presented [30], comprising two convolutional layers, two average pooling layers, and one fully connected layer.
+Which is the first concept of the LeNet architecture.
+In 1998, the LeNet-5 model was presented [32], comprising three convolutional layers, two pooling layers, and two fully connected layers.
+It was designed to efficiently recognize handwritten postal codes, outperforming traditional methods like the multilayer-perceptron model.
+CNNs gained significant traction in the 2010s, primarily due to their excellent performance on large datasets like ImageNet [15], which includes millions of images across 1,000 classes.
+A typical CNN consists of an input layer, multiple hidden layers, and an output layer.
 The input layer receives the image and forwards its pixel values into the network.
@@
 The hidden layers, situated between the input and output layers, extract relevant features using convolution, activation, and pooling operations.
-Each hidden layer processes information from the input or preceding layers, enabling iterative learning.
-This information is then relayed to subsequent layers, culminating in the final prediction or classification at the output layer.
+Each hidden layer receives information from either the input or preceding layers, facilitating iterative learning.
+This information is then passed to subsequent layers, culminating in the final prediction or classification at the output layer.
 Kiranyaz et al.
 [28] introduced a 1D CNN for disease-specific ECG classification.
-A 1D CNN, tailored for one-dimensional data, is particularly suitable for signal data, time-series data, and text data.
-In this architecture, both the input and the convolution filter are one-dimensional; the filter slides over the input in the convolution layer to perform operations.
+A 1D CNN is a variant of CNNs tailored for one-dimensional data, making it particularly suitable for signal data, time-series data, and text data.
+In this architecture, both the input and the convolution filter are one-dimensional; the filter slides over the input in the convolution layer to execute operations.
 II-B Homomorphic Encryption (HE)
@@
 In UniHENN, we use the CKKS scheme, which is advantageous for encrypting vectors of real or complex numbers.
-Before encrypting with CKKS, it is crucial to create a plaintext structure that mimics the input data.
-The number of slots specifies the vector size that can be encrypted and is set during parameter selection.
-CKKS supports three basic operations: addition, multiplication of ciphertexts, and rotation of ciphertext.
-Let $N$ be the total degree of the CKKS parameter then the number of slots is $N/2$.
-Two real vectors $\mathbf{v_{1}},\mathbf{v_{2}}$ and $C(\mathbf{v_{1}}),C(\mathbf{v_{2}})$ denotes the ciphertext of the vector $\mathbf{v_{1}},\mathbf{v_{2}}$.
-(P) means that the operation is defined between a ciphertext and a plaintext, and (C) means that the operation is defined between two ciphertexts.
+CKKS encryption requires structuring plaintext to mirror input data.
+The number of slots, defined during parameter selection, determines the encryptable vector size.
+CKKS supports three fundamental ciphertext operations: addition, multiplication, and rotation.
+Let $N$ be the total degree of the CKKS parameter, then the number of slots is $N/2$.
+Two real vectors $\mathbf{v_{1}},\mathbf{v_{2}}$ and $C(\mathbf{v_{1}}),C(\mathbf{v_{2}})$ denote the ciphertext of the vectors $\mathbf{v_{1}},\mathbf{v_{2}}$.
+(P) indicates that the operation is defined between a ciphertext and a plaintext, and (C) indicates that the operation is defined between two ciphertexts.
 Then the operations, addition ($Add$), multiplication ($Mul$), and rotation $(Rot)$ can be represented as follows:
@@
   • Multiplication (C): $Mul(C(\mathbf{v_{1}}),C(\mathbf{v_{2}}))=C(\mathbf{v_{1}}\times\mathbf{v_{2}})$
-  • Rotation :
-    $Rot(C(\mathbf{v}),r)=C(v_{r},v_{r+1},...,v_{N/2-1},v_{0},...,v_{r-1})$
-    where $\mathbf{v}=(v_{1},v_{2},...,v_{N/2})$ and $r$ is a positive integer.
-The notation +,×+$\times$+ , × in the plaintext addition and multiplication are elementwisely defined.
-The depth specifies the maximum number of multiplications, which affects the noise level in the ciphertext, and exceeding this limit may compromise decryption.
-The level specifies the allowed number of the multiplication.
-The depth is fixed when the parameter is determined but the level can be decreased when multiplications are conducted.
-For practical HE implementations, various libraries are available, including SEAL-Python [2], Lattigo [39], HElib [18], and OpenFHE [5].
+  • Rotation: $,Rot(C(\mathbf{v}),r)=C(v_{r},v_{r+1},\ldots,v_{N/2-1},v_{0},$$\ldots,v_{r-1})$, where $\mathbf{v}=(v_{0},v_{1},\ldots,v_{N/2-1})$ and $r$ is a positive integer.
+The operations $+$ and $\times$ represent elementwise addition and multiplication in the plaintext space, respectively.
+The depth parameter specifies the maximum number of sequential multiplications that can be performed while ensuring correct decryption, as each multiplication operation increases ciphertext noise.
+Exceeding this limit may result in decryption failure.
+While depth is predetermined during parameter selection, level dynamically indicates the remaining multiplication capacity, decreasing with each multiplication operation executed.
+For practical HE implementations, several general-purpose HE libraries are available, including SEAL-Python [2], Lattigo [40], HElib [20], and OpenFHE [5].
 We selected SEAL-Python for its efficient support of the CKKS scheme.
-II-C HE Libraries
-Several libraries have been developed in the HE field to facilitate practical implementations. Among these, SEAL-Python [44], Lattigo [39], HElib [1], and OpenFHE [5] are popularly used.
-SEAL [44] supports various HE schemes (BFV and CKKS) and incorporates optimizations to speed up homomorphic operations, such as ciphertext packing and batching.
-For UniHENN, we chose SEAL-Python, a Python adaptation of the original C++ SEAL library.
-We selected SEAL-Python for its user-friendly interface.
-However, UniHENN could also be implemented with other libraries.
-Lattigo [39], written in Go, focuses on ring-learning-with-errors-based HE primitives and multiparty-HE-based secure protocols.
-Its cross-platform compatibility and optimized protocols enable fast and secure computation.
-HElib [1] specializes in the BGV and CKKS schemes, incorporating bootstrapping and several optimizations to speed up homomorphic evaluation.
-It effectively uses the Smart-Vercauteren ciphertext packing techniques and the Gentry-Halevi-Smart optimization.
-Also, it provides an assembly language for HE, which includes low-level routines, multithreading capabilities, and more.
-OpenFHE [5], supporting both bootstrapping and various hardware acceleration backends, utilizes a standard hardware abstraction layer to ensure compatibility and efficiency.
+Building on this foundation, we developed UniHENN, a HE-based framework specifically optimized for CNN inference.
+While currently implemented with SEAL-Python, UniHENN is designed to be adaptable to other libraries.
 III Overview of UniHENN
@@
   3. 3.
-    UniHENN uses average pooling, eliminating multiplication.
+    UniHENN opts for average pooling, which eliminates the need for multiplication, thereby reducing operation time.
 This configuration allows for more layers within given parameter settings, giving service providers more flexibility to incorporate average pooling without concern for operation count.
@@
 These encrypted images are consolidated into a single ciphertext and sent to a cloud service specializing in data analytics.
-Each service (i.e., Service 1, Service 2, …, Service N) uses its unique CNN model for data processing, executing specific algorithms and computations on the received ciphertext with the evaluation key.
+Each service (i.e., Service 1, Service 2, …, Service $K$) uses its unique CNN model for data processing, executing specific algorithms and computations on the received ciphertext with the evaluation key.
 Each CNN model has a distinct layer architecture and optimized parameters.
@@
 This framework allows the client to benefit from multiple CNN models while preserving the confidentiality of the data, indicating a significant step forward in privacy-preserving machine learning.
-In Table I, we summarize the important notations used in this paper with UniHENN.
+Table I summarizes the notations used throughout this paper.
 TABLE I: Notations used for UniHENN.
-Notation
-Definition
-$N$
-Number of slots of ciphertext in power of 2.
-$D$
-Depth (or level) of ciphertext, representing the current limit on the number of multiplication operations allowed for the ciphertext.
-$CH_{in},CH_{out}$
-Number of input and output channels in a layer.
-$DAT_{in},DAT_{out}$
-Number of input and output data in a fully connected (FC) layer.
-$R_{q}$
-Polynomial quotient ring $(\mathbb{Z}_{q}/\langle X^{N}+1\rangle)$, used to create a vector space for the ciphertext.
-$W_{img},H_{img}$
-Width and height of the used image data.
-$K$
-Width and height of the kernel.
-$W_{in},H_{in}$
-Width and height of input data for this layer.
-$W_{out},H_{out}$
-Width and height of output data for this layer.
-$S_{total}$
-Used in Section IV-E. The value is the product of each convolutional layer’s stride and the kernel sizes of all average pooling layers.
+Notation Definition
+$N$ Degree of polynomial. It is in the form of a power of two.
+$D,L$ Depth and level of ciphertext. This represents the current remaining limit on the number of multiplication operations allowed for the ciphertext.
+$CH_{in},CH_{out}$ Number of input and output channels in layer
+$DAT_{in},DAT_{out}$ Number of input and output data in FC layer
+$R_{q}$ Polynomial quotient ring $(\mathbb{Z}_{q}/\langle X^{N}+1\rangle)$. It is used to create a vector space of the ciphertext.
+$W_{img},H_{img}$ Width and height of the input image data
+$K$ Width and height of the kernel
+$W_{in},H_{in}$ Width and height of input data from this layer
+$W_{out},H_{out}$ Width and height of output data from this layer
+$S_{total}$ It is used in Section IV-E. The value is the multiplication of each convolutional layer’s stride and the kernel sizes of all average pooling layers.
 III-A Construction of Input Data
@@
 The remaining space will be used in the batch operations.
-Figure 2 illustrates an example of data transformation for encryption.
+Figure 2 shows an example of data transformation for encryption.
 The input data consists of $W_{img}\times H_{img}$ numbers.
 We flatten the original data row by row, as shown in Figure 2.
-Implementing this flattening procedure ensures universal compatibility of the input data with ML models, enhancing its adaptability to the algorithms used by ML services.
+This flattening procedure ensures universal compatibility of the input data with ML models, enhancing adaptability to algorithms used by ML services.
 Figure 2: Data transform for UniHENN.
@@
 By examining the given CNN model structure, we can pre-calculate this size in advance, thereby preventing the overlap of input data results.
-For example, as illustrated in Figure 3, if 800 slots are needed to process the input data through the CNN model, then the $i$-th input data would be rotated to the right by $(i-1)\times 800$ slots, and the rotated vector is added to the ciphertext.
+Figure 3 illustrates one example.
+In this case, 800 slots are needed to process the input data through the CNN model.
+Thus, the $i$-th input data would be rotated to the right by $(i-1)\times 800$ slots, and the rotated vector is added to the ciphertext.
 UniHENN can integrate any number of ciphertexts into a single one, provided the total size of the encrypted data does not exceed the number of slots.
@@
 So, the CKKS parameter is determined and encrypted before the CNN model is defined for reusability of the input ciphertext independent of the model.
-But, the multiplication and rotation time of HE is linearly increased dependent on the depth.
+But, the multiplication and rotation time of HE is increased dependent on the level.
 Thus, the best choice of the CKKS parameter is determined by the number and structure of the CNN model’s layers.
 However, in this scenario, the depth is defined before when the CNN model is fixed, which means that the depth can be larger than the exact number of total levels of the CNN’s model.
-Thus, we suggest the Drop-DEPTH, a novel approach to reduce the depth that is optimized when the target CNN model.
-If the input ciphertext has depth $D$ and the target CNN model’s level is $L$ where $D\leq L$ then the Drop-DEPTH reduces the depth of the input ciphertext as $L$ with multiply $D-L$ ciphertexts that are the encryption of vectors where all elements are consists of $1$.
-By these simple structure, time of execution is very fast (In our experiments in Section VII, the time of Drop-DEPTH operation does not exceed $100ms$.
+Thus, we suggest the Drop-Level, a novel approach to reduce the level that is optimized when the target CNN model.
+If the input ciphertext has depth $D$ and the target CNN model’s level is $L$ where $L\leq D$ then the Drop-Level reduces the level of the input ciphertext as $L$ with multiply $D-L$ ciphertexts that are the encryption of vectors where all elements are consists of $1$.
+By these simple structures, the time of execution is very fast (In our experiments in Section VII, the time of Drop-Level operation does not exceed $100ms$.
 IV-B Construction of the Convolutional Layer
-The im2col encoding is a highly efficient algorithm for converting multidimensional data into matrix form to facilitate convolution operations.
+The im2col encoding is a popular and efficient algorithm for transforming multidimensional data into matrix form to facilitate convolution operations.
 However, it requires precise arrangement of input matrix elements, which is challenging to achieve on ciphertexts.
-Therefore, TenSEAL [7], a sophisticated open-source HE-enabled ML SDK, performs this operation in a preprocessing step before encryption, allowing only a single convolution.
-To overcome the limitations of im2col, we introduce a new method for computing convolutional layers that does not rely on this encoding.
-Our approach uses flattened input data, eliminating the need for a specific arrangement of input matrix elements and thereby increasing flexibility.
-In convolutional layers, a kernel traverses an input image with each movement termed a stride.
-The kernel executes a convolution operation at each stride, multiplying its values with corresponding input image values and summing the results.
-The input image dimensions are denoted as $(W_{img},H_{img})$, and the stride as $S$.
-And for simplicity, we define the width and height are equal and denoted as $K$.
-The total number of kernel movements is given by Equation 1:
-$\left(\frac{W_{img}-S}{K}\right)\left(\frac{H_{img}-S}{K}\right)$ (1)
-The total number of operations executed by a convolutional layer is influenced by both the size of the input image and the kernel, as indicated by Equation 1.
-However, given that the input image is usually significantly larger than the kernel, we can approximate that the total number of operations primarily depends on the kernel size.
+Therefore, TenSEAL [7], a sophisticated open-source HE-enabled ML SDK, performs this operation in a preprocessing step before encryption, allowing only a single convolutional layer.
+To simplify this process and improve flexibility, we propose a new method for constructing convolutional layers without relying on im2col.
+Our approach uses flattened input data, eliminating the need for a specific arrangement of input matrix elements as follows:
+  1. 1.
+    Kernel and Stride: In convolutional layers, a small matrix called a kernel moves over the input image.
+Each movement of the kernel is called a stride.
+At each position, the kernel multiplies its values with the corresponding values in the input image and sums the results.
+  2. 2.
+    Input and Kernel Dimensions: The dimensions of the input image are denoted as $(W_{img},H_{img})$, with a stride of $S$.
+For simplicity, we assume the kernel has equal width and height, both represented by $K$.
+  3. 3.
+    Flattening the Input: In the context of HE, we flatten the two-dimensional input data into a one-dimensional form. This simplifies the convolution process under encryption.
+  4. 4.
+    Rotation and Multiplication: The convolution operation involves rotating the input data and multiplying it with the kernel.
+The number of rotations equals the number of elements in the kernel.
+Each element in the input data is multiplied by the corresponding element in the kernel, and the results are summed up.
+  5. 5.
+    Output: The output of the convolution is an array with values spaced according to the stride interval.
+Figures 4 and 5 illustrate the convolution operations on two-dimensional and flattened input data, respectively.
 Figure 4: Convolution operations on the two-dimensional input data and kernel.
-As shown in Figure 4, the input data dimension for a basic convolution is two-dimensional.
-However, input data must be restricted to one-dimensional form in the context of HE.
-The input data is flattened, as demonstrated in Figure 5, using the method outlined in Section III-A.
-For simplicity, we assume that both the input data and the kernel are two-dimensional squares and disregard bias operations.
-Figure 5: Convolution operations on the flattened input data.
-In Figure 5, the number of rotation operations equals the number of elements in the kernel.
-The input data is incrementally increased by the original kernel size $K$, followed by a leftward rotation operation after each increment.
-After adding $K\times W_{img}$ cipertexts, another leftward rotation is performed, incrementing by one at each step.
-The detailed algorithm of these mechanism is described in Algorithm 1.
-All rotated ciphertexts are then multiplied by the kernel vector, which is specially structured to contain a single kernel element with all other positions set to zero.
-This eliminates unrelated elements in the ciphertext during convolution, as depicted in Figure 6.
-For 1D convolution, the logic remains the same, except that the one-dimensional data consists of a single row.
-The output is an array spaced by the stride interval, as shown in Figure 7.
-Figure 6: Flattening a two-dimensional kernel into a one-dimensional array.
-Figure 7: Output of convolution operations on one-dimensional data.
-This approach has the advantage of decoupling the number of operations from the size of the image.
-Instead, it performs convolutional operations on transformed two-dimensional data and limits rotations to the square of the kernel size $K$.
-Previous approaches have used im2col encoding [7] or performed operations on individual data values [16].
-However, the im2col method restricts the model to a single convolutional layer and requires model-specific encoding.
-Methods that operate on each data value are considerably slower.
+Figure 5: Convolution operations on the flattened data.
+Figure 5 illustrates the convolution process: the input data is incrementally shifted by $S$, $K$ times, with leftward rotations after each shift.
+These rotated ciphertexts are then multiplied by a specially structured kernel vector containing a single kernel element, effectively eliminating unrelated elements during convolution.
+Algorithm 1 provides a detailed description of this convolution layer construction.
 Algorithm 1 Construction of the convolution layer
@@
 - Interval : $I_{out}$
-Procedure: *(All ciphertexts and operation are in $\mathcal{R}_{q}^{2}$)
+Procedure: *(All ciphertexts and operations are in $\mathcal{R}_{q}^{2}$)
 for $i=0$ to $CH_{in}-1$ do
-     for $p=0$ to $K-1$ do
-         for $q=0$ to $K-1$ do
-              $C_{rot(i,p,q)}$
-              $\leftarrow Rot\left(C_{in(i)},I_{in}\times(q+W_{img}\times p)\right)$
+     for $j=0$ to $K-1$ do
+         for $k=0$ to $K-1$ do
+              $C_{rot(i,j,k)}$
+              $\leftarrow Rot\left(C_{in(i)},I_{in}\times(k+W_{img}\times j)\right)$
          end for
@@
 for $o=0$ to $CH_{out}-1$ do
-     $C_{out(o)}$
-     $\leftarrow\sum_{i=0}^{CH_{in}-1}\sum_{p=0}^{K-1}\sum_{q=0}^{K-1}Mul\left(C_{rot(i,p,q)},K_{(o,i,p,q)}\right)$
+     $C_{out(t)}$
+     $\leftarrow\sum_{i=0}^{CH_{in}-1}\sum_{j=0}^{K-1}\sum_{k=0}^{K-1}Mul\left(C_{rot(i,j,k)},K_{(o,i,j,k)}\right)$
      $C_{out(o)}\leftarrow Add(C_{out(o)},B_{(o)})$
@@
 return $C_{out},I_{out}$
-The number of multiplications in the convolution layer of UniHENN is $\mathcal{O}(CH_{in}\cdot CH_{out}\cdot K^{2})$, and the number of rotations in the convolution layer of UniHENN is $\mathcal{O}(CH_{in}\cdot K^{2})$.
-As stated in  [32], the complexity of a single multiplication is $\mathcal{O}(N\cdot D)$, and the complexity of a single rotation is $\mathcal{O}(N\cdot\log N\cdot D^{2})$.
-Therefore, the complexity of the convolutional layer of UniHENN is $\mathcal{O}(N\cdot D\cdot CH_{in}\cdot K^{2}\cdot(CH_{out}+\log N\cdot D))$.
+This approach improves the convolution process by decoupling the number of operations from the image size and limiting rotations to the square of the kernel size $K$.
+The number of plaintext multiplications in the convolution layer of our method is $\mathcal{O}(CH_{in}\cdot CH_{out}\cdot K^{2})$ and the number of rotations is $\mathcal{O}(CH_{in}\cdot K^{2})$.
+Considering the complexity of plaintext multiplication $\mathcal{O}(N\cdot L)$ and rotation $\mathcal{O}(N\cdot\log N\cdot L^{2})$ as stated in [33], the overall complexity of the convolutional layer is $\mathcal{O}(N\cdot L\cdot CH_{in}\cdot K^{2}\cdot(CH_{out}+\log N\cdot L))$.
 IV-C Construction of the Average Pooling Layer
-CNNs use pooling layers to reduce the input size of convolutional layers.
-The most commonly used pooling layers are max pooling, min pooling, and average pooling.
-Max and min pooling require many multiplications [34], while average pooling requires one.
-Our architecture aims to provide a CNN inference with high speed, we only consider the CNN models without bootstrapping.
-But, the max pooling and min pooling consume lots of multiplications so it is needed to use the bootstrapping with these pooling layers.
-Therefore, in UniHENN, we adopt average pooling instead of max pooling and min pooling.
-Originally, the average operation is a multiplication of the constant $1/c^{2}$, where $c$ is the size of the kernel.
-However, in UniHENN, we modify the average pooling layer to eliminate the need for constant multiplication.
-Our main idea is to apply the constant multiplication in one of the convolutional layers or flatten layers instead of the average pooling layer.
-It provides the same functionality as the general average pooling but reduces multiplication only changing the order of multiplication of $1/c^{2}$ in the previous layer.
-So, it can choose lower depth, resulting in efficiency in the overall operation.
-The choice is determined by the layer that follows average pooling.
-If an activation layer follows, the constant multiplication is passed to either the flatten layer or the next convolutional layer, as this layer is linear.
-Using the linearity $h(x)=Ax+b$, we find $h(x/c^{2})=A(x/c^{2})+b=(1/c^{2})Ax+b$, allowing us to use the weight $(1/c^{2})A$ instead of $A$.
+CNNs use pooling layers to reduce input size, with max, min, and average pooling being common.
+Our architecture prioritizes high-speed CNN inference without bootstrapping.
+Ablation studies on a LeNet-5-like model revealed comparable accuracies for min (0.989), max (0.992), and average (0.985) pooling.
+In HE, max and min pooling require numerous multiplications [35], necessitating costly bootstrapping, while average pooling needs only one multiplication.
+Given the negligible performance difference and significant computational savings, we implemented average pooling in UniHENN.
+We optimize the average pooling layer by eliminating the constant multiplication by $1/c^{2}$ (where $c$ is the kernel size).
+Instead, we apply this multiplication in a preceding convolutional or flatten layer.
+This approach maintains functionality while reducing the number of multiplications and lowering depth, enhancing overall efficiency.
+The implementation depends on the subsequent layer.
+If followed by an activation layer, the constant multiplication is moved to the preceding flatten or convolutional layer.
+For a linear layer $h(x)=Ax+b$, we use $h(x/c^{2})=(1/c^{2})Ax+b$, incorporating the scaling factor into the weight matrix.
+This strategy preserves the average pooling effect while optimizing computational resources.
 When an activation function such as the square function or the approximate ReLU function follows average pooling, we cannot apply the above logic directly because these functions are not linear.
@@
 $(W_{img}+1)\times(c-1)$ (1)
+Figure 6: Flattening a two-dimensional kernel into a one-dimensional array.
 IV-D Construction of the Fully Connected (FC) Layer
@@
 Next, append the segments of each array with a negative index to the end of the array.
-Finally, as depicted in Figure 9 (c), divide the matrix lengthwise by the output size and concatenate the parts horizontally; the weight matrix $M_{w}$ then becomes $M_{rot}$ as shown in Figure 9 (d).
-Note that generating an optimized weight matrix (from Figure 9 (a) to Figure 9 (d)) involves modifying the weights in plaintext.
-Therefore, the operation time is negligible.
-Algorithm 2 provides the details of the process for constructing the fully connected layer.
+Finally, as depicted in Figure 8 (c), divide the matrix lengthwise by the output size and concatenate the parts horizontally; the weight matrix $M_{w}$ then becomes as shown in Figure 8 (d).
+Note that generating an optimized weight matrix (from Figure 8 (a) to Figure 8 (d)) involves modifying the weights in plaintext; however, the operation time is negligible.
 In the inference process, the input ciphertext is rotated and multiplied by the weight matrix $M_{w}$, as depicted in Figure 8 (e).
@@
 - Ciphertext: $C_{out}\in\mathcal{R}_{q}^{2}$
+Procedure: *(All ciphertexts and operations are in $\mathcal{R}_{q}^{2}$)
+$M_{rot}\leftarrow$ Apply $M_{w}$ from step 1 to 4 in Figure 8
+$W_{vec}=\lceil DAT_{in}/DAT_{out}\rceil\times DAT_{out}$
+$C_{sum}\leftarrow C(0)$
 for $o=0$ to $CH_{out}-1$ do
-     $V_{1}\leftarrow M_{rot(o)}[:DAT_{out}-o]+[0]\times o$
-     $C_{out(o)}\leftarrow Mul(Rot(C_{in},o),V_{1})$
-     $V_{2}\leftarrow[0]\times(DAT_{out}-o)+M_{rot(o)}[DAT_{out}-o:]$
-     $V_{3}\leftarrow Mul(Rot(C_{in},o-DAT_{out}),V_{2})$
-     $C_{out(o)}\leftarrow Add(C_{out(o)},V_{3})$
+     $V_{1}\leftarrow M_{rot(o)}[:W_{vec}-o]+[0]\times o$
+     $C_{sum}\leftarrow Add(C_{sum},Mul(Rot(C_{in},o),V_{1}))$
+     $V_{2}\leftarrow[0]\times(W_{vec}-o)+M_{rot(o)}[W_{vec}-o:]$
+     $C_{sum}\leftarrow Add(C_{sum},Mul(Rot(C_{in},o-W_{vec}),V_{2}))$
 end for
-for $i=1$ to $\lceil DAT_{in}/DAT_{out}\rceil-1$ do
-     $C_{out}\leftarrow Add(C_{out},Rot(C_{out},i\times DAT_{in}))$
+$C_{out}\leftarrow C(0)$
+for $i=0$ to $\lceil DAT_{in}/DAT_{out}\rceil-1$ do
+     $C_{out}\leftarrow Add(C_{out},Rot(C_{sum},i\times DAT_{out}))$
 end for
-$C_{out}\leftarrow Add(C_{out},B)$
-return $C_{out}$
-The number of multiplications in the fully connected (FC) layer of UniHENN is $\mathcal{O}(DAT_{out})$, and the number of rotations in the FC layer of UniHENN is $\mathcal{O}(DAT_{in}/DAT_{out}+DAT_{out})$.
-This implies that if $DAT_{in}$ is smaller than the square of $DAT_{out}$, the rotations can be denoted as $\mathcal{O}(DAT_{out})$.
-In this scenario, the time complexity of the FC layer depends solely on the output size of the FC layer.
-Thus, the complexity of the FC layer of UniHENN is $\mathcal{O}((DAT_{in}/DAT_{out}+DAT_{out})\cdot N\cdot\log N\cdot D^{2})$.
+$C_{out}\leftarrow Add(C_{out},B)$ return $C_{out}$
+The number of plaintext multiplications of the FC layer in UniHENN is $\mathcal{O}(DAT_{out})$, and the number of rotations is $\mathcal{O}(DAT_{in}/DAT_{out}+DAT_{out})$.
+This means that if $DAT_{in}$ is smaller than the square of $DAT_{out}$, it can be denoted as $\mathcal{O}(DAT_{out})$.
+In this case, the time complexity of the FC layer depends only on the output size.
+Therefore, the complexity of the FC layer in UniHENN is $\mathcal{O}((DAT_{in}/DAT_{out}+DAT_{out})\cdot N\cdot\log N\cdot L^{2})$.
 IV-E Construction of the Flatten Layer
@@
 Figure 9: Removing row interval when the previous layer is either a convolutional layer or an approximate ReLU layer.
-After passing through the convolution layer or fully connected layer, the valid data in the output ciphertext will have offset, which we call row interval.
-We propose an algorithm to remove the row interval for two situations: the preceding layer is an average pooling layer or the preceding layer is not an average pooling.
+After passing through the convolution layer or fully connected layer, the valid data in the output ciphertext will be offset, which we call row interval.
+We propose an algorithm to remove the row interval for two situations: whether the preceding layer is an average pooling layer or not.
 As detailed in Section IV-C, the square layer does not handle constant multiplication, so it is not considered when identifying the previous layer.
-As illustrated in Figure 10, if the previous layer is an average pooling layer, the given input ciphertexts have invalid values in the intervals between each set of values.
-Furthermore, since we deferred the constant multiplication ($1/c^{2}$) in the pooling layer, we need to multiply $1/c^{2}$ to the ciphertexts.
-Section IV-E2 describes the three steps to address this issue.
-If the preceding layer is either a convolutional layer or an approximate ReLU layer, it will contain some zero values between each set of non-zero values.
+If the previous layer is either a convolutional layer or an approximate ReLU layer, it may contain some zero values between each set of non-zero values.
 We can exploit this to efficiently execute the flatten operation by stacking the non-zero values using rotation and addition.
@@
 IV-E2 Removing Invalid Data and Row Interval
+If the previous layer is an average pooling layer, invalid values exist in the intervals between each set of valid data points.
+Moreover, the required constant multiplication (assuming a factor of $1/c^{2}$) has not been applied during the average pooling operation.
+Consequently, we must eliminate these invalid data points and perform the necessary multiplication.
+To address this issue, we suggest a three-step process as follows.
 Figure 10: Removing invalid data between valid data when the previous layer is an average pooling layer.
@@
 Thus, only a single multiplication operation would be needed.
-The number of multiplications in the flatten layer of UniHENN is $\mathcal{O}(W_{in}+H_{in})$, and the number of rotations in the flatten layer of UniHENN is $\mathcal{O}(CH_{in}\cdot S_{total}+W_{in}+H_{in})$.
-Therefore the complexity of flatten layer of UniHENN is $\mathcal{O}((CH_{in}\cdot S_{total}+W_{in}+H_{in})\cdot N\cdot\log N\cdot D^{2})$
+IV-E5 Time Complexity
+The number of plaintext multiplications in the flatten layer of UniHENN is $\mathcal{O}(W_{in}+H_{in})$ and the number of rotations in the flatten layer of UniHENN is $\mathcal{O}(CH_{in}\cdot S_{total}+W_{in}+H_{in})$.
+Therefore the complexity of flatten layer of UniHENN is $\mathcal{O}((CH_{in}\cdot S_{total}+W_{in}+H_{in})\cdot N\cdot\log N\cdot L^{2})$
 IV-F Construction of the Activation Layer
@@
 Theorem 1.
-Let $W_{img}\times H_{img}$ be the image size and $L$ the number of total convolutional layers.
-Denote $(W_{in(i)},H_{in(i)})$ and $(W_{out(i)},H_{out(i)})$ as the input and output sizes of the $i$-th convolutional layer, and the kernel size is $(KW_{i},KH_{i})$, stride size is $(SW_{i},SH_{i})$, and padding size is $(PW_{i},PH_{i})$ for all $i\in\left[1,L\right]$.
+Let $W_{img}\times H_{img}$ be the image size and $N_{layer}$ the number of total convolutional layers.
+Denote $(W_{in(i)},H_{in(i)})$ and $(W_{out(i)},H_{out(i)})$ as the input and output sizes of the $i$-th convolutional layer, and the kernel size is $(W_{ker(i)},H_{ker(i)})$, stride size is $(W_{st(i)},H_{st(i)})$, and padding size is $(W_{pad(i)},H_{pad(i)})$ for all $i\in\left[1,N_{layer}\right]$.
 Additionally, let
-$SH_{(i)}=\prod\limits_{j=1}^{i}SH_{j}$ (3)
-Then, $H_{img}\geq H_{out(i)}\times SH_{(i)}$, where $KH_{i}\geq SH_{i}\geq 1$ for all $i\in\left[1,L\right]$.
+$H_{[st(i)]}=\prod\limits_{j=1}^{i}H_{st(j)}$ (2)
+Then, $H_{img}\geq H_{out(i)}\times H_{[st(i)]}$, where $H_{ker(i)}\geq H_{st(i)}\geq 1$ for all $i\in\left[1,N_{layer}\right]$.
 Proof.
 We obtain that $H_{out(i)}$ :
-$H_{out(i)}=\left[\frac{H_{in(i)}-KH_{i}}{SH_{i}}\right]+1$ (4)
+$H_{out(i)}=\left[\frac{H_{in(i)}-H_{ker(i)}}{H_{st(i)}}\right]+1$ (3)
 Given the hypothesis, the following inequality holds:
-$H_{out(i)}\times SH_{i}$ $=\left(\left[\frac{H_{in(i)}-KH_{i}}{SH_{i}}\right]+1\right)\times SH_{i}$ (5)
-$=\left[\frac{H_{in(i)}-KH_{i}+SH_{i}}{SH_{i}}\right]\times SH_{i}$ (6)
-$\leq\left[\frac{H_{in(i)}}{SH_{i}}\right]\times SH_{i}\leq H_{in(i)}=H_{out(i-1)}$ (7)
+$H_{out(i)}\times H_{st(i)}$ $=\left(\left[\frac{H_{in(i)}-H_{ker(i)}}{H_{st(i)}}\right]+1\right)\times H_{st(i)}$ (4)
+$=\left[\frac{H_{in(i)}-H_{ker(i)}+H_{st(i)}}{H_{st(i)}}\right]\times H_{st(i)}$ (5)
+$\leq\left[\frac{H_{in(i)}}{H_{st(i)}}\right]\times H_{st(i)}\leq H_{in(i)}=H_{out(i-1)}$ (6)
 When $i>1$, $H_{in(i)}=H_{out(i-1)}$. As proven by mathematical induction, inserting an average pooling layer or an activation function between convolutional layers does not change the size. ∎
-Note that after passing through the $i$-th convolutional layer, the interval in each column of data is $W_{img}\times(SH_{(i)}-1)$ (detailed in  IV-B).
+Note that after passing through the $i$-th convolutional layer, the interval in each column of data is $W_{img}\times(H_{[st(i)]}-1)$ (detailed in  IV-B).
 Given $W_{img}$ columns of data or fewer, the image size of $H_{img}\times W_{img}$ is not exceeded by Theorem 1.
@@
 Denote
-$SW_{(i)}=\prod\limits_{j=1}^{i}SW_{j}$ (8)
-Then, $W_{img}\geq W_{out(i)}\times SW_{(i)}$, where $KW_{i}\geq SW_{i}\geq 1$ for all $i\in\left[1,L\right]$.
+$W_{[st(i)]}=\prod\limits_{j=1}^{i}W_{st(j)}$ (7)
+Then, $W_{img}\geq W_{out(i)}\times W_{[st(i)]}$, where $W_{ker(i)}\geq W_{st(i)}\geq 1$ for all $i\in\left[1,N_{layer}\right]$.
 Proof.
@@
 According to Theorem 1, the number of slots in the ciphertext data remains unchanged when passing through UniHENN’s average pooling.
-However, to prevent invalid values from occupying the positions of other used data, an additional space of $(W_{img}+1)\times(c-1)$ is required (detailed in Section IV-C).
+However, to prevent invalid values from occupying the positions of other used data, an additional space of $(W_{img}+1)\times(c-1)$ is required (detailed in IV-C).
 V-C Flatten Layer
@@
 VI-A Experimental Settings
-We used a server service on the NAVER Cloud platform [42] with the following specifications: 16 vCPU cores (Intel(R) Xeon(R) Gold 5220 CPU @ 2.20GHz), 64GB of memory, and a 50GB SSD.
+We used a server on the NAVER Cloud platform [44] with the following specifications: 16 vCPU cores (Intel(R) Xeon(R) Gold 5220 CPU @ 2.20GHz), 64GB of memory, and a 50GB SSD.
 VI-B Datasets
-We utilize four different datasets for each model: MNIST [15], CIFAR-10 [26], USPS [19], and ECG [38].
+We utilized four different datasets for our models: MNIST [16], CIFAR-10 [29], USPS [21], and ECG [39].
 VI-B1 MNIST Dataset
-MNIST is a dataset [15] consisting of 70,000 grayscale images of handwritten digits, each with a size of $28\times 28$ pixels.
-The dataset contains 10 classes, representing the digits from 0 to 9.
-The dataset consists of two parts: 60,000 data for training and 10,000 data for testing.
-We use these settings in our experiments.
+The MNIST dataset [16] consists of 70,000 grayscale images of handwritten digits, each sized $28\times 28$ pixels, with 10 classes representing the digits from 0 to 9.
+We used 60,000 images for training and 10,000 images for testing.
 VI-B2 CIFAR-10 Dataset
-CIFAR-10 is a dataset [26] comprising 60,000 RGB images, each with dimensions $3\times 32\times 32$ pixels.
-Each image is labeled with one of 10 classes.
-The dataset consists of two parts: 60,000 data for training and 10,000 data for testing.
-We use these settings in our experiments.
+The CIFAR-10 dataset [29] includes 60,000 RGB images, each sized $32\times 32$ pixels, with 10 classes. We used 50,000 images for training and 10,000 images for testing.
 VI-B3 USPS Dataset
-The USPS dataset [19] consists of 9,298 grayscale images of handwritten digits, each with dimensions $16\times 16$ pixels.
-The dataset also contains 10 classes, representing the digits from 0 to 9.
-The dataset consists of two parts: 7,291 data for training and 2,007 data for testing.
-We use these settings in our experiments.
+The USPS dataset [21] contains 9,298 grayscale images of handwritten digits, each sized $16\times 16$ pixels, with 10 classes. We used 7,291 images for training and 2,007 images for testing.
 VI-B4 ECG Dataset
-The ECG dataset from the MIT-BIH arrhythmia database [38] is used for the performance evaluation of our proposed patient-specific ECG approach.
-It has a total of 109,446 samples with a sampling frequency of 125Hz.
-The signals represent electrocardiogram (ECG) shapes of heartbeats in normal cases and cases affected by various arrhythmias and myocardial infarction.
-Each data point is labeled with one of 5 classes.
-In Abuadbba et al.
-[3], they pre-processed the ECG dataset to choose a total of 43 out of 48 records are selected and each signal was rescaled by the min-max normalization.
-And, choose the biorthogonal wavelet as decomposition and reconstruction wavelet function.
-We use 13,245 data for training and 13,245 data for testing which is in the https://github.com/SharifAbuadbba/split-learning-1D.
-We use a preprocessed dataset generated by Abuadbba et al. [3], where each data point is one-dimensional and has 128 columns.
-VI-B5 Hyperparameter Settings for Training
-We trained all models with the learning rate of 0.001 and the batch size of 32. We trained the MNIST and USPS models for 15 epochs, the CIFAR-10 model for 25 epochs, and the ECG model for 30 epochs.
-VI-C Library and Parameter Setup
-We used SEAL-Python [2], a library that provides Python APIs for performing homomorphic operations like key generation, encryption, decryption, addition, multiplication, and rotation.
-The reason for using the SEAL-based library is that baseline models TenSEAL, and PyCrCNN in our experiments, are based on SEAL.
-So, we use it to balance the experiment.
-And, we use the pytorch and torchvision library for training.
-The detailed information of libraries and the experimental setup is described in the GitHub site : https://github.com/hm-choi/uni-henn.
-The specific parameters used in our experiments are detailed in Table II.
-All security parameters meet the 128-bit security level as described in [43].
+The ECG dataset from the MIT-BIH arrhythmia database [39] consists of 109,446 samples at a sampling frequency of 125 Hz.
+It includes signals representing ECG shapes of heartbeats in normal cases and various arrhythmias.
+Following the preprocessing steps by Abuadbba et al.
+[4], we used a subset of 26,490 samples, with 13,245 samples for training and 13,245 samples for testing, available at https://github.com/SharifAbuadbba/split-learning-1D.
+VI-C Hyperparameter Settings for Training
+We trained all models with a learning rate of 0.001 and a batch size of 32.
+The MNIST and USPS models were trained for 15 epochs, the CIFAR-10 model for 25 epochs, and the ECG model for 30 epochs, based on their convergence behavior.
+VI-D Library and Parameter Setup
+We used SEAL-Python [2] for homomorphic operations, ensuring consistency with our baseline models TenSEAL and PyCrCNN, which are also SEAL-based.
+PyTorch and torchvision were used for model training.
+Detailed information about libraries and the experimental setup is available at our GitHub repository: https://github.com/hm-choi/uni-henn.
+Table II provides the specific parameters used in our experiments.
+All security parameters comply with the 128-bit security level as described in [45].
 TABLE II: Parameters used in UniHENN.
-$PK$ and $SK$ represent the size of the public key and the secret key.
-$GK$ represents the size of the Galois key, $RK$ represents the size of the set of Relinearization keys, and # mult means the total number of allowed multiplications.
+$PK$: public key size (encryption key), $SK$: secret key size (decryption key), $GK$: Galois key size, $RK$: Relinearization keys size.
+$GK$ and $RK$ together form the evaluation key.
+# mult: total allowed multiplications.
 # slots 8,192
@@
 # mult (depth) 11
-The public key and secret key are also referred to as the encryption key and decryption key, respectively.
-The set consisting of the Galois key and the relinearization keys is also known as the evaluation key.
 VII Experimental Results
@@
 This algorithm allows TenSEAL to achieve highly optimized and efficient inference times for convolution operations.
-However, this efficiency comes at the cost of flexibility: TenSEAL’s architecture only supports models with a single convolutional layer.
-Therefore, TenSEAL is not a viable option for CNN architectures with multiple convolutional layers.
-PyCrCNN is an alternative to TenSEAL that does not employ the im2col algorithm, making it applicable to a broader range of CNN architectures.
-However, this flexibility comes with a significant drawback: PyCrCNN exhibits slower inference times than TenSEAL.
-Our experiments show that UniHENN strikes a balance between these extremes.
-It achieves inference times comparable to those of TenSEAL while also supporting CNN architectures with multiple convolutional layers.
-This makes UniHENN a more versatile solution for HE-based deep learning inference.
+However, this efficiency comes at the cost of flexibility: TenSEAL’s architecture only supports models with a single convolutional layer, making it unsuitable for CNN architectures with multiple convolutional layers.
+PyCrCNN, on the other hand, does not employ the im2col algorithm, making it applicable to a broader range of CNN architectures.
+However, this flexibility results in slower inference times compared to TenSEAL.
+Our experiments show that UniHENN successfully balances both flexibility and efficiency.
+It achieves inference times comparable to those of TenSEAL while supporting CNN architectures with multiple convolutional layers, making UniHENN a more versatile solution for HE-based deep learning inference.
   • Adaptability of UniHENN for Various CNN Model Architectures.
@@
 Our successful implementation of LeNet-5 in UniHENN highlights the system’s ability to handle large and complex CNN models effectively.
-We also consider a 1D CNN architecture particularly useful for sequence-based tasks such as time-series analysis, natural language processing, and certain bioinformatics applications.
+We also considered a 1D CNN architecture particularly useful for sequence-based tasks such as time-series analysis, natural language processing, and certain bioinformatics applications.
 This capability sets UniHENN apart from solutions like TenSEAL, which is constrained to supporting only single convolutional layer models and does not offer support for 1D CNNs.
@@
 Table III provides detailed specifications for this model architecture.
-The Conv2D parameters–$ch_{in}$, $ch_{out}$, size, and stride–represent the number of input channels, the number of output channels, kernel size, and stride for a two-dimensional convolutional layer, respectively.
-Similarly, $data_{in}$ and $data_{out}$ in the FC layer indicate the input and output dimensions.
+The Conv2D parameters – $CH_{in}$, $CH_{out}$, $K$, and $S$ – represent the number of input channels, the number of output channels, kernel size, and stride for a two-dimensional convolutional layer, respectively.
+Similarly, $DAT_{in}$ and $DAT_{out}$ in the FC layer indicate the input and output dimensions.
 We use these notations to represent all other remaining models presented in this paper.
@@
 Layer Parameter # Mult
-Conv2d $ch_{in}$=1, $ch_{out}$=8, size=4, stride=3 1
+Conv2d $CH_{in}$ = 1, $CH_{out}$ = 8, $K$ = 4, $S$ = 3 1
 Square - 1
 Flatten - 2
-FC1 $data_{in}$ = 648, $data_{out}$ = 64 1
+FC1 $DAT_{in}$ = 648, $DAT_{out}$ = 64 1
 Square - 1
-FC2 $data_{in}$ = 64, $data_{out}$ = 10 1
+FC2 $DAT_{in}$ = 64, $DAT_{out}$ = 10 1
 Total #Mult - 7
@@
 TenSEAL achieved this efficiency by optimizing convolution operations on the input data using the im2col algorithm.
-UniHENN took 16.248 seconds, making it 2.58 times slower than TenSEAL.
+UniHENN took 16.247 seconds, making it 2.6 times slower than TenSEAL.
 PyCrCNN was the slowest, requiring 154.494 seconds.
@@
 TenSEAL and PyCrCNN do not offer parallel processing for multiple samples, resulting in total times that increase proportionally with the number of input samples.
-To process ten input samples together, UniHENN still takes 16.248 seconds, approximately 3.9 times and 94.6 times faster than TenSEAL’s 63.706 seconds and PyCrCNN’s 1537.228 seconds, respectively.
+To process ten input samples together, UniHENN still takes 16.247 seconds, approximately 3.9 times and 94.6 times faster than TenSEAL’s 63.706 seconds and PyCrCNN’s 1537.227 seconds, respectively.
 TABLE IV: Comparison of the average (with standard deviation) inference time in seconds on the MNIST dataset between PyCrCNN and UniHENN for the $M_{1}$ model architecture.
@@
 # of samples 1 10 1 10 1 & 10
-Drop DEPTH - - - -
+Drop Level - - - -
 0.065
@@
 154.494
-1537.228
-16.248
+1537.227
+16.247
 We conducted additional experiments to determine the number of samples at which UniHENN starts to outperform TenSEAL.
@@
 In this experiment, we utilize the LeNet-1 model, denoted as $M_{2}$.
-The model accuracy of $M_{2}$ is 98.62%, slightly higher than $M_{1}$.
+The accuracy of $M_{2}$ is 98.62%, slightly higher than $M_{1}$.
 The detailed specifications of these model architectures can be found in Table V.
-Implementing the hyperbolic tangent function (tanh) in HE is computationally challenging; therefore, we have substituted the activation function from the tanh to the square function.
+Implementing the hyperbolic tangent function (tanh) in HE is computationally challenging; therefore, we have substituted the activation function from tanh to the square function.
 We find that the modified LeNet-1 model with the square activation achieves an accuracy of 98.62% comparable to the 98.41% accuracy of the original LeNet-1 model when tested on a 10,000 sample MNIST dataset.
@@
 Layer Parameter #Mult
-Conv2d $ch_{in}$=1, $ch_{out}$=4, size=5, stride=1 1
+Conv2d $CH_{in}$ = 1, $CH_{out}$ = 4, $K$ = 5, $S$ = 1 1
 Square - 1
 AvgPool2d kernel size = 2 0
-Conv2d $ch_{in}$=4, $ch_{out}$=12, size=5, stride=1 1
+Conv2d $CH_{in}$ = 4, $CH_{out}$ = 12, $K$ = 5, $S$ = 1 1
 Square - 1
@@
 Flatten - 2
-FC1 $data_{in}$=192, $data_{out}$=10 1
+FC1 $DAT_{in}$ = 192, $DAT_{out}$ = 10 1
 Total #Mult - 7
@@
 Model PyCrCNN UniHENN
-Drop DEPTH - 0.066 (0.003)
+Drop Level - 0.066 (0.003)
 Conv2d 169.182 (0.891) 3.715 (0.048)
@@
 FC1 6.527 (0.070) 0.522 (0.014)
-Total 794.064 (2.279) 30.090 (0.155)
-In Table VI, UniHENN takes 30.090 seconds, making it approximately 26.4 times faster than PyCrCNN, which takes 794.064 seconds.
+Total 800.591 (2.279) 30.089 (0.155)
+In Table VI, UniHENN takes 30.089 seconds, making it approximately 26.6 times faster than PyCrCNN, which takes 800.591 seconds.
 In this experiment, most of the computational time for both UniHENN and PyCrCNN is consumed in the second convolutional layer.
@@
 Layer Parameter #Mult
-Conv2d $ch_{in}$=1, $ch_{out}$=6, size=3, stride=1 1
+Conv2d $CH_{in}$ = 1, $CH_{out}$ = 6, $K$ = 3, $S$ = 1 1
 Approx ReLU $f(x)=0.375373+0.5x+0.117071x^{2}$ 2
@@
 Flatten - 2
-FC1 $data_{in}$ = 1014, $data_{out}$ = 120 1
+FC1 $DAT_{in}$ = 1014, $DAT_{out}$ = 120 1
 Approx ReLU $f(x)=0.375373+0.5x+0.117071x^{2}$ 2
-FC2 $data_{in}$ = 120, $data_{out}$ = 10 1
+FC2 $DAT_{in}$ = 120, $DAT_{out}$ = 10 1
 Total #Mult - 9
@@
 The Flatten and FC1 layers are the most time-consuming, taking an average of 9.603 seconds and 15.557 seconds, respectively.
-These two layers alone contribute significantly to the total inference time of 29.104 seconds.
+These two layers alone contribute significantly to the total inference time of 29.105 seconds.
 While UniHENN is significantly slower than non-HE models, it provides enhanced privacy and security, which may be crucial for specific applications or compliance requirements.
-The inference time of 29.104 seconds is still under 30 seconds, suggesting that UniHENN is practical for real-world applications, especially in contexts where data security is paramount.
+The inference time of 29.105 seconds is still under 30 seconds, suggesting that UniHENN is practical for real-world applications, especially in contexts where data security is paramount.
 This inference time could be considered acceptable depending on the specific use case and the sensitivity of the data being processed.
@@
 Model UniHENN
-Drop DEPTH 0.035 (0.001)
+Drop Level 0.035 (0.001)
 Conv2d 1.863 (0.029)
@@
 FC2 0.459 (0.010)
-Total 29.104 (0.171)
+Total 29.105 (0.171)
 VII-D Experiment 4: Adaptibility of UniHENN for LeNet-5
@@
 Layer Parameter #Mult
-Conv2d $ch_{in}$=1, $ch_{out}$=6, size=5, stride=1 1
+Conv2d $CH_{in}$ = 1, $CH_{out}$ = 6, $K$ = 5, $S$ = 1 1
 Square - 1
 AvgPool2d kernel size = 2 0
-Conv2d $ch_{in}$=6, $ch_{out}$=16, size=5, stride=1 1
+Conv2d $CH_{in}$ = 6, $CH_{out}$ = 16, $K$ = 5, $S$ = 1 1
 Square - 1
 AvgPool2d kernel size = 2 0
-Conv2d $ch_{in}$=16, $ch_{out}$=120, size=5, stride=1 1
+Conv2d $CH_{in}$ = 16, $CH_{out}$ = 120, $K$ = 5, $S$ = 1 1
 Square - 1
 Flatten - 0
-FC1 $data_{in}=120$, $data_{out}=84$ 1
+FC1 $DAT_{in}$ = 120, $DAT_{out}$ = 84 1
 Square - 1
-FC2 $data_{in}=84$, $data_{out}=10$ 1
+FC2 $DAT_{in}$ = 84, $DAT_{out}$ = 10 1
 Total #Mult - 9
 Table X shows a layer-by-layer breakdown of the inference time for the $M_{4}$ model architecture.
-The results show that the inference time for $M_{4}$ is 740.127 seconds.
+The results show that the inference time for $M_{4}$ is 740.128 seconds.
 This is substantially slower than the previous experiments, which could be attributed to the model’s increased complexity.
@@
 Layer UniHENN
-Drop DEPTH 0.035 (0.002)
+Drop Level 0.035 (0.002)
 Conv2d 5.244 (0.048)
@@
 FC2 0.454 (0.009)
-Total 740.127 (1.381)
+Total 740.128 (1.381)
 VII-E Experiment 5: Adaptability of UniHENN for CNN Models on Color Images
@@
 Layer Parameter #Mult
-Conv2d $ch_{in}$=3, $ch_{out}$=16, size=3, stride=1 1
+Conv2d $CH_{in}$ = 3, $CH_{out}$ = 16, $K$ = 3, $S$ = 1 1
 Square - 1
 AvgPool2d kernel size = 2 0
-Conv2d $ch_{in}$=16, $ch_{out}$=64, size=4, stride=1 1
+Conv2d $CH_{in}$ = 16, $CH_{out}$ = 64, $K$ = 4, $S$ = 1 1
 Square - 1
 AvgPool2d kernel size = 2 0
-Conv2d $ch_{in}$=64, $ch_{out}$=128, size=3, stride=1 1
+Conv2d $CH_{in}$ = 64, $CH_{out}$ = 128, $K$ = 3, $S$ = 1 1
 Square - 1
@@
 Flatten - 1
-FC1 $data_{in}=128$, $data_{out}=10$ 1
+FC1 $DAT_{in}$ = 128, $DAT_{out}$ = 10 1
 Total #Mult - 8
-Table XII presents the experimental results, showing that the total inference time for UniHENN is approximately 20 minutes.
+Table XII presents the experimental results, showing that the total inference time for UniHENN is approximately 21 minutes.
 While this may seem long, it is important to note that the convolutional layers are particularly time-consuming, requiring about 256.000 seconds for the first layer and 938.446 seconds for the second.
@@
 Layer UniHENN
-Drop DEPTH 0.155 (0.004)
+Drop Level 0.155 (0.004)
 Conv2d 8.935 (0.062)
 Square 0.703 (0.011)
-AvgPool2d 1.891 (0.022)
+AvgPool2d 0.429 (0.004)
 Conv2d 256.000 (0.867)
 Square 1.839 (0.023)
-AvgPool2d 4.466 (0.049)
+AvgPool2d 1.891 (0.022)
 Conv2d 938.446 (1.884)
@@
 Total 1251.444 (2.532)
-VII-F Experiment 6: Comparison of Inference Time of UniHENN on Grayscale Images
+VII-F Experiment 6: Evaluation of Inference Time of UniHENN on Grayscale Images
 To assess the adaptability of UniHENN to diverse datasets, we performed experiments using the USPS dataset and the $M_{6}$ model architecture.
@@
 Layer Parameter #Mult
-Conv2d $ch_{in}$=1, $ch_{out}$=6, size=4, stride=2 1
+Conv2d $CH_{in}$ = 1, $CH_{out}$ = 6, $K$ = 4, $S$ = 2 1
 Square - 1
 Flatten - 2
-FC1 $data_{in}$ = 294, $data_{out}$ = 64 1
+FC1 $DAT_{in}$ = 294, $DAT_{out}$ = 64 1
 Square - 1
-FC2 $data_{in}$ = 64, $data_{out}$ = 10 1
+FC2 $DAT_{in}$ = 64, $DAT_{out}$ = 10 1
 Total #Mult - 7
-Table XIV shows that UniHENN is approximately 6 times faster than PyCrCNN, with a total inference time of 12.307 seconds compared to PyCrCNN’s 71.139 seconds.
+Table XIV shows that UniHENN is approximately 6 times faster than PyCrCNN, with a total inference time of 12.309 seconds compared to PyCrCNN’s 71.139 seconds.
 Interestingly, the FC1 layer in PyCrCNN consumes a significant portion of the time (42.792 seconds), likely due to its large input size of 294.
@@
 Layer PyCrCNN UniHENN
-Drop DEPTH - 0.066 (0.002)
+Drop Level - 0.066 (0.002)
 Conv2d 14.441 (0.125) 2.662 (0.029)
@@
 FC2 1.166 (0.015) 0.384 (0.009)
-Total 71.139 (0.198) 12.307 (0.092)
-VII-G Experiment 7: Comparison of Inference Time for a 1D CNN Model between UniHENN and PyCrCNN
+Total 71.139 (0.198) 12.309 (0.092)
+VII-G Experiment 7: Evaluation of Inference Time of UniHENN for a 1D CNN Model
 To evaluate the adaptability of UniHENN for 1D CNN models, we implemented a 1D CNN model, denoted as $M_{7}$, for processing ECG data.
@@
 Layer Parameter #Mult
-Conv1d $ch_{in}$=1, $ch_{out}$=2, size=2, stride=2 1
+Conv1d $CH_{in}$ = 1, $CH_{out}$ = 2, $K$ = 2, $S$ = 2 1
 Square - 1
-Conv1d $ch_{in}$=2, $ch_{out}$=4, size=2, stride=2 1
+Conv1d $CH_{in}$ = 2, $CH_{out}$ = 4, $K$ = 2, $S$ = 2 1
 Flatten - 1
-FC1 $data_{in}$ = 128, $data_{out}$ = 32 1
+FC1 $DAT_{in}$ = 128, $DAT_{out}$ = 32 1
 Square - 1
-FC2 $data_{in}$ = 32, $data_{out}$ = 5 1
+FC2 $DAT_{in}$ = 32, $DAT_{out}$ = 5 1
 Total #Mult - 7
-Table XVI shows that UniHENN is approximately 3.03 times faster than PyCrCNN, recording a total inference time of 5.119 seconds compared to PyCrCNN’s 15.513 seconds.
+Table XVI shows that UniHENN is approximately 3.0 times faster than PyCrCNN, recording a total inference time of 5.119 seconds compared to PyCrCNN’s 15.514 seconds.
 This indicates that UniHENN is also more efficient even for relatively smaller models, such as 1D CNNs, where the computational time for convolutional layers is not as extensive.
@@
 Layer PyCrCNN UniHENN
-Drop DEPTH - 0.069 (0.007)
+Drop Level - 0.069 (0.007)
 Conv1d 0.556 (0.011) 0.113 (0.003)
@@
 FC2 0.242 (0.006) 0.180 (0.016)
-Total 15.513 (0.114) 5.119 (0.039)
+Total 15.514 (0.114) 5.119 (0.039)
 VII-H Validation of inference results on encrypted data
@@
 Therefore, it is crucial to validate UniHENN’s inference results by comparing them with the outcomes of the original plaintext inference.
-We selected $M_{4}$ and $M_{5}$ as representative models for validation, as the errors were minimal in other models.
+We selected $M_{4}$ and $M_{5}$ as representative models for validation, as the errors were very small in the case of other models.
 We evaluated the $M_{4}$ model with 2,000 samples from the MNIST dataset and the $M_{5}$ model with 2,000 samples from the CIFAR-10 dataset.
@@
 VII-I Impact of HE Parameters
-We conducted an additional analysis to investigate how the inference time and encrypted inference error vary depending on the parameters of the CKKS scheme, the HE algorithm we employ.
+We conducted an additional analysis to investigate how the inference time and encrypted inference error vary depending on the parameters of the CKKS scheme.
 The CKKS scheme’s operation time and decrypted result accuracy are influenced by several parameters, specifically # slots, scale, and depth.
@@
 To address this, we set the depth as high as possible and then optimize by fine-tuning the input ciphertext.
-Figure 15: Average inference time (seconds) for $M_{1}$ with depth.
+Figure 14: Average inference time for $M_{1}$ with depth.
 VII-I2 Impact of Scale on Error
@@
 Figure 15: Error in $M_{1}$ with scale.
+VII-J Comparison of UniHENN and DP-based CNN
+This section compares our UniHENN framework with differential privacy (DP) for privacy-preserving CNNs, providing insights into the trade-offs between HE and DP-based approaches.
+We implemented DP-SGD [3] using Opacus [52].
+We conducted experiments on the MNIST dataset with model $M_{4}$, comprising 60,000 training and 10,000 test images.
+The Adam optimizer was used with a learning rate of 0.001, training for 15 epochs with a batch size of 32.
+For the DP-SGD implementation, we set the maximum per gradient norm to 1.0 and used $\delta=0.00001$, satisfying the condition $\delta<1/60,000$ for the MNIST training set size.
+We varied $\epsilon$ from 5 to 0.1 during the 15-epoch training.
+While there is no definitive guideline for choosing $\epsilon$, NIST’s post [41] suggests that $\epsilon$ between 0 and 5 provides strong privacy protection.
+Thus, our selected $\epsilon$ values with $\delta=0.00001$ would represent suitable candidates for DP-SGD-based training.
+Tables XVII and XVIII present the accuracy and memory consumption results, respectively.
+TABLE XVII: Accuracy comparison between UniHENN (U) and DP-based CNN (D) for various privacy budgets ($\epsilon$) (30 runs, standard deviations in parentheses).
+Framework U D ($\epsilon$=5) D ($\epsilon$=1) D ($\epsilon$=0.5) D ($\epsilon$=0.1)
+Accuracy (%) 98.91 84.19 (2.59) 84.57 (2.29) 83.73 (1.98) 80.63 (1.54)
+TABLE XVIII: Memory consumption comparison during inference (30 runs, standard deviations in parentheses).
+Framework U D
+Memory (MiB) 3677.64 (1.05) 3.91 (0.10)
+Our results reveal significant trade-offs between HE-based and DP-based privacy-preserving techniques in CNNs.
+UniHENN achieves superior accuracy (98.91%) compared to the DP-based CNN, which ranges from 84.57% to 80.63% as $\epsilon$ decreases.
+This demonstrates UniHENN’s ability to maintain model performance across varying privacy parameters.
+However, this high accuracy comes at a substantial computational cost: UniHENN consumes about 941 times more memory (3677.64 MiB) than the DP-based approach (3.91 MiB) due to the intensive nature of HE operations.
+These findings highlight the critical resource trade-offs in privacy-preserving machine learning, emphasizing the need to balance accuracy and computational efficiency when selecting privacy-preserving techniques for specific applications.
+The choice between HE and DP approaches will depend on the particular use case, with UniHENN offering superior accuracy at the cost of higher computational resources, while DP-based methods provide a more memory-efficient solution with a trade-off in accuracy.
 VIII Limitations
-We propose an optimized CNN model inference mechanism based on HE, UniHENN. UniHENN uses batching to reduce inference time, but it has three limitations.
-As deep learning has evolved, it has been used in many applications beyond image classification [48], such as detecting attacks on the web [45] or analyzing resource extraction [47].
-These DNN models require a large number of operations, especially multiplication operations.
-To apply homomorphic encryption, multiplication operations must be used within a certain number of times depending on the parameter values used, or bootstrapping must be used.
-Due to the nature of deep learning models with a large number of multiplication operations, bootstrapping is required, but bootstrapping has the limitation of slowing down the computation speed significantly.
-UniHENN does not support multi-core architecture or GPUs.
-As we tried to deploy our source codes, we chose to use SEAL-Python, an open-source homomorphic encryption framework.
-But most open-source homomorphic encryption frameworks, including SEAL-Python, are suitable for only CPU and single-thread architecture.
-Nowadays, some studies for optimizing HE operations in the GPU environment [6, 40] are conducted.
-The main idea of the study is to improve performance of the HE-operation using the parallelism design.
-So, to address these limitations, we plan to extend our architecture that support the multi-core environment and GPU environment in the software side with the previous works for future work.
-Since UniHENN utilizes the batch system to reduce computational time.
-UniHENN will be efficient when analyzing sensitive data at scale or when clients infer multiple data at once.
-However, UniHENN may not be efficient when it is needed to infer single data because the design point of UniHENN is focused on batch operations.
-To overcome this limitation, we plan to improve the computational efficiency that the operations constituting the convolutional layer or FC layer operation will be distributed in the space where the batch operation was performed to increase efficiency.
+We propose an optimized CNN model inference mechanism based on HE, UniHENN. While UniHENN uses batching to reduce inference time, it has three key limitations.
+Firstly, complex deep learning models often require a large number of operations, particularly multiplications.
+In HE, multiplication operations must be limited based on parameter values, or bootstrapping must be used.
+Due to the extensive multiplication operations in deep learning models, bootstrapping is necessary, significantly slowing computation speed.
+Secondly, UniHENN does not currently support multi-core architectures or GPUs.
+Our deployment uses SEAL-Python, an open-source HE framework suited for CPU and single-thread architectures.
+Most open-source HE frameworks, including SEAL-Python, are designed this way.
+Recent studies have optimized HE operations for GPUs [6, 42], focusing on parallelism to enhance performance.
+To address these limitations, we plan to extend our architecture to support multi-core and GPU environments.
+Our future work will involve optimizing key HE operations for parallel execution on GPUs, implementing efficient memory management techniques, and ensuring proper synchronization mechanisms to maximize computational performance.
+Finally, UniHENN is designed to reduce computational time through batching, making it efficient for analyzing sensitive data at scale or when inferring multiple data points simultaneously.
+However, due to its focus on batch operations, UniHENN may be inefficient for inferring single data points.
+To overcome this, we plan to improve computational efficiency by distributing operations within the convolutional and FC layers in the space where batch operations are performed, thereby increasing overall efficiency.
 IX Related Work
-Dowlin et al.
-[17] proposed CryptoNets, the first HE-based CNN inference scheme.
-It utilizes square activation instead of ReLU to reduce the total computation cost with a low number of multiplications.
-However, the input ciphertext size and weights of CryptoNets are very large because they encrypt each pixel of the input image and each element of the weights individually.
-Chillotti et al.
-[12] introduced an enhanced version of CryptoNets, applying low-degree polynomials for functions like ReLU, Softplus, and Swish and incorporating techniques like pruning and quantization for more efficient homomorphic operations.
+IX-A Libraries of HE
+Several prominent HE libraries are available, including SEAL [46], Lattigo [40], HElib [1], and OpenFHE [5].
+Each offers unique features: SEAL [46] supports BFV and CKKS schemes with optimizations like ciphertext packing.
+We used SEAL-Python, a Python port of the C++ implementation, for our work.
+Lattigo [40] implements Ring-LWE-based HE primitives and Multiparty-HE protocols in Go, enabling cross-platform builds and optimized secure computation.
+HElib [1] supports BGV and CKKS schemes with bootstrapping, focusing on Smart-Vercauteren ciphertext packing and Gentry-Halevi-Smart optimization.
+It also provides an HE assembly language for low-level control.
+OpenFHE [5] offers bootstrapping and hardware acceleration using a standard Hardware Abstraction Layer (HAL).
+IX-B Machine learning implementation with HE
+To provide privacy-preserving ML services, many previous studies implement HE-friendly neural networks.
+They focus on two challenges: “How to build HE-friendly model architecture, especially, the activation layer” and “How to build an efficient HE-based CNN (or DNN) system.”
+In the first case, there were efforts to replace activation functions with polynomial functions.
+CryptoNets [19] use a square function to replace the ReLU activation function.
+Chou et al.
+[13] proposed some techniques to replace activation functions such as ReLU, Softplus, and Swish with low-degree polynomials and use pruning and quantization for the efficiency of homomorphic operations.
 Ishiyama et al.
-[20] adopted an approximation of Google’s Swish activation function using two and four-degree polynomials, coupled with batch normalization, aiming to minimize the accuracy gap between traditional and polynomial-based activation functions.
-In 2021, Benaissa et al.
-presented TenSEAL [7], a prominent open-source tool that supports HE-based CNN inference.
-It utilizes the im2col method to efficiently compute the convolutional layer.
-However, it cannot support multiple convolutional layers and does not support batch operation.
-In this paper, we propose UniHENN to address these issues.
-Aharoni et al.
-[4] presented an encryption data structure, the tile tensor, for efficiently calculating deep neural networks containing CNNs.
-It splits and encrypts the input data considering the model structure.
-However, the encryption of input data in this scheme depends on the model architecture; sharing the client’s ciphertext with a third party, such as a public cloud service, is challenging because the CNN model’s structure is usually undisclosed.
-Lee et al.
-[32] proposed a novel low-latency model optimization strategy using an enhanced HerPN block and masking convolution, significantly reducing multiplication operations in the convolution layer.
-However, their approach does not consider the design of FC layers.
-In this paper, we present how to implement all necessary layers for CNNs in HE efficiently.
+[22] adopt Google’s Swish activation function with two- and four-degree polynomials with batch normalization to reduce the errors between Swish and approximated polynomials.
+However, replacing activation functions with low-degree polynomials reduces model accuracy and is only applicable to shallow models.
+To overcome these limitations, Park et al.
+[43] proposed the HerPN block, which can replace the batch normalization and ReLU block by utilizing the Hermite polynomial.
+Recently, Lee et al.
+[33] proposed PP-DNN, a low-latency model optimization solution focusing on convolution and approximate ReLU operations without bootstrapping, supporting CNN models like ResNet-34 with HE.
+In contrast, UniHENN offers comprehensive optimizations across all layer types, including convolutional, pooling, fully connected, and flatten layers, resulting in more efficient operations overall.
+Additionally, UniHENN introduces a batch system for multi-input inference and optimizes element sizes for batch operations, which are not addressed in PP-DNN.
+In the second case, many studies [23, 43, 18] interact with clients to compute non-linear operations.
+The server sends the encrypted intermediate results to the client, and the client computes the non-linear functions with decrypted intermediate results.
+In this way, the model accuracy can be preserved, but the communication overhead is increased.
 X Conclusion
-In this paper, we introduce a privacy-preserving machine learning framework, UniHENN, addressing a significant challenge in cloud computing.
-UniHENN enables the construction of HE-enabled CNN models without relying on the traditional im2col operation, enhancing compatibility with diverse ML models and reinforcing privacy in cloud-based machine learning.
-Our extensive evaluation of UniHENN across four public datasets using seven different CNNs, including six 2D CNNs and a 1D CNN, demonstrates its robustness and efficacy.
-Notably, UniHENN maintains accuracy levels comparable to unencrypted models, highlighting its practical applicability without performance compromise.
-A key feature of UniHENN is its batch mechanism support, significantly accelerating processing times for multiple datasets concurrently, a breakthrough for cloud computing.
-Our findings indicate that a CNN constructed with UniHENN outperforms the model $M1$ (comprising one convolutional layer and two fully connected layers) by almost 3.921 times using the TenSEAL library, demonstrating a substantial improvement in processing speed, crucial for real-world applications requiring rapid data processing.
-We publicly release UniHENN’s source code at https://github.com/hm-choi/uni-henn to enhance its impact, promoting widespread use.
-In the future, we aim to optimize UniHENN for multi-thread architectures, leveraging the additional space available in batch computations.
-This optimization will involve dividing the operations of each layer to enable simultaneous processing, thereby enhancing the efficiency of UniHENN.
-We aim to establish UniHENN as a practical tool for secure and efficient cloud computing in machine learning applications.
+[…]